### v2

# Speech Robust Bench: A Robustness Benchmark For Speech Recognition

Abstract
As Automatic Speech Recognition (ASR) models become ever more pervasive, it is important to ensure that they make reliable predictions under corruptions present in the physical and digital world. We propose Speech Robust Bench (SRB), a comprehensive benchmark for evaluating the robustness of ASR models to diverse corruptions. SRB is composed of 114 input perturbations which simulate an heterogeneous range of corruptions that ASR models may encounter when deployed in the wild. We use SRB to evaluate the robustness of several state-of-the-art ASR models and observe that model size and certain modeling choices such as the use of discrete representations, or self-training appear to be conducive to robustness. We extend this analysis to measure the robustness of ASR models on data from various demographic subgroups, namely English and Spanish speakers, and males and females. Our results revealed noticeable disparities in the model’s robustness across subgroups. We believe that SRB will significantly facilitate future research towards robust ASR models, by making it easier to conduct comprehensive and comparable robustness evaluations.
1 Introduction
As novel ML models continue to be developed and deployed at an ever-increasing rate, it has become crucial to ensure that their robustness to corruptions arising from a myriad of sources, including the environment, sensing apparatus, and even malicious actors is well understood. To this end, prior works have developed comprehensive robustness benchmarks, particularly for vision [1, 2, 3, 4] and natural language processing models [5, 6], that evaluate several aspects of a model’s performance under a variety of input perturbations. These benchmarks have proven to be invaluable to the advancement of research into more robust models because (1) they unify robustness evaluations, thus enabling meaningful comparisons across models and allowing progress to be accurately tracked, and (2) they make it easier for researchers to comprehensively evaluate the robustness of their models by aggregating a diverse and representative set of perturbations in a single benchmark.
While several robustness benchmark datasets exist for Automatic Speech Recognition (ASR) models [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], none of the currently existing ones is in any sense comprehensive: each benchmark measures the model’s robustness w.r.t. to one or a few specific types of perturbations, thereby putting the onus on model developers to find and collect all the relevant benchmarks to evaluate their model comprehensively. This has often resulted in model developers evaluating their models on disparate benchmarks  [19, 20, 21, 22], which makes it hard to reliably compare performance and robustness across models. Recently, Huggingface Open ASR Leaderboard [23] has sought to unify ASR model evaluations by developing a benchmark consisting of several real-world speech datasets. Although evaluating models on exclusively natural data may accurately reflect their real-world performance, it is generally not informative about the specific types of perturbations the models are weak or strong against, because the noise sources present in these datasets are not controlled or even fully known. For example, the crowdsourced recordings in Common Voice [24] contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers. Additionally, crucial digital perturbations like special effects, computer-generated speech, and adversarial examples, to which humans are usually invariant, are largely overlooked by existing benchmarks.
In this paper, we propose Speech Robust Bench (SRB), a benchmark for comprehensively evaluating the robustness of ASR models to input perturbations and corruptions. SRB is designed to address the aforementioned major shortcomings of existing ASR robustness benchmarks, i.e., that (1) they are often specialized and thus are not individually comprehensive, (2) even taken together, they overlook important perturbation types. SRB addresses these shortcomings by curating a large comprehensive bank of 114 perturbations and inherently noisy data sources. The perturbation bank includes distortions and variations caused by the environment, recording equipment, special effects, computer-generated speech, and significantly, adversarial attacks, which are often overlooked by existing benchmarks. When selecting inherently noisy data sources, we take special care to ensure that the sources of variation in each data source are limited and known, so that the models’ specific weaknesses can be identified. These sources contain speech with non-native English accents and inter-personal conversations, giving rise to articulatory and lexical variability in speech.
To highlight the need for and the benefits of doing systematic and comprehensive robustness assessment, we evaluate the robustness of several popular ASR models (§4.1) using SRB. We observe that Whisper [19] is the most robust on average among the models we tested, even outperforming the more recent Canary [25]. We conduct further analyses to disentangle the effects of model and training data size, revealing that larger models tend to be more robust than smaller models, even if the latter are trained on significantly more data. We further extend our analysis by evaluating the models’ robustness for the various population sub-groups, namely, English and non-English (Spanish) speakers, and male and female speakers. We find that significant disparities exist across these sub-groups, thus identifying issues with fairness of the models and highlighting areas where future work could provide clear improvements. Besides pinpointing robustness issues, this demonstrates the utility of SRB for fairness evaluations that consider robustness disparities across models as well.
To summarize we make the following contributions:
  • We present SRB, a comprehensive robustness benchmark for ASR models, enabling direct and easy comparison of robustness evaluations between models to facilitate progress.
  • We demonstrate the use of SRB by conducting a fine-grained robustness analysis for several popular models. We extend our analysis by using SRB to uncover disparities in the robustness of ASR for various sub-groups of speakers. This highlights the broad utility of this benchmark to the field of trustworthy AI.
  • To facilitate out-of-the-box robustness evaluations for the research community, we have publicly released a large dataset (https://huggingface.co/datasets/mshah1/speech_robust_bench) containing perturbed versions of LibriSpeech [26], TEDLIUM [27] and Spanish Multilingual LibriSpeech [28] test sets, and accented speech from common voice and segemented near- and far-field audios from CHiME-6 [10] and AMI[8].
  • We open source our code with clear documentation to allow for easy reproducibility and extensibility.(https://github.com/ahmedshah1494/speech_robust_bench)
2 Related Work
2.1 Robust Automatic Speech Recognition
Several techniques have been proposed for making Automatic Speech Recognition (ASR) models robust to input perturbations, such as noise and other signal corruptions [29]. We can divide these techniques into two high-level categories: i) model-based and ii) feature-based. Model-based techniques modify the models to make them more robust. Examples of such approaches include adapting pre-trained models [30, 31], de-noising the audio before processing [32, 33], and training ASR models on noisy data [22]. Since model-based strategies generally require access to noisy data [29], they are most effective if the sources of noise, and/or the exact environment in which the ASR model will be deployed in are known, and one can gather data to represent them. Feature-based approaches, on the other hand, involve developing handcrafted features, invariant to noise and corruptions in the signal [29]. Several of these features are inspired by biological audition [34, 35, 36], while others use signal processing techniques [29]. Generally, these methods are designed to extract the components of the audio signal salient for speech production and perception, while discarding irrelevant components [37]. Consequently, they do not require precise knowledge of the environment and noise distributions. Recently, however, handcrafted features have fallen out of favor, and have been replaced by features learned via end-to-end training of deep learning models on large amounts of data [38, 39, 22, 19]. Proponents of these techniques posit that models trained on larger datasets become more robust. Our evaluations in § 4 reveal that there are several input perturbations against which smaller models trained on less data outperform larger models trained on more data.
2.2 Adversarial Robustness
Adversarial perturbations can change the response of a model when added to their inputs, but are either imperceptible to humans or perceptually and semantically irrelevant enough to be ignored by them [40, 41]. Adversarially perturbed inputs are known as adversarial attacks. They can be targeted (aiming to change a prediction to a specific incorrect class), or un-targeted (aiming to change a prediction to any incorrect class [42]). The design of adversarial attacks is determined by the level of knowledge the attacker is assumed to have about the target model. Attacks that assume full knowledge of the target model’s architecture and weights (white-box threat model) often use gradient-based optimization techniques [40, 41, 43, 44, 42]. Attackers who do not have any knowledge of the target model’s architecture and only have query access to it (black-box threat model) typically use gradient-free optimization methods [45, 46, 47, 48, 49, 50]. An intriguing property of adversarial perturbations is that they transfer between models [51], and inputs [42, 52], i.e. perturbations designed for one model/input may be effective against others as well. Our SRB includes two types of white box adversarial attacks; those that generate perturbations: 1) specific to each input [43], 2) that cause models to mis-transcribe multiple inputs [52].
2.3 Robustness Benchmarks for Speech
Several robustness benchmark datasets exist for Automatic Speech Recognition (ASR) models [7, 8, 9, 10, 11, 12, 9, 13, 14, 15, 16, 17, 18, 53, 54, 55]. Unfortunately, these benchmarks do not comprehensively evaluate model robustness. While past works such as [54] did indeed propose benchmarks containing diverse perturbations, many of the recently proposed benchmarks measure the model’s robustness to one or a few specific types of perturbations. For example, some datasets [15, 17, 18, 16] focus exclusively on reverberant speech, while others focus only on multi-speaker scenarios [8, 7], environmental noise [14, 10, 9, 53], or accented speech [56, 57]. In this situation, the onus is on model developers to collect all relevant benchmarks and comprehensively evaluate their model. Furthermore, certain types of digital perturbations, like special effects, computer-generated speech, and adversarial examples, which humans are usually invariant to (and thus ASR models are expected to be as well) are largely overlooked by existing benchmarks. To facilitate future research, SRB combines several benchmark datasets containing known noise sources to allow researchers to comprehensively evaluate the robustness of their models using a single benchmark.
In the absence of comprehensive benchmarks, prior works have tried to evaluate ASR model robustness in various ways. For example, several works have used subsets of existing benchmarks to evaluate the model robustness [19, 20, 21], while others [22, 19, 58] have evaluated the robustness of models by computing their transcription accuracy on multiple natural speech datasets. A key shortcoming of these approaches is that the choice of benchmarks and datasets varies across studies, making it difficult to compare results. Recently, Huggingface Open ASR Leaderboard [23] has sought to mitigate this issue of inconsistent evaluation setups by developing a benchmark comprising several real-world speech datasets. Although evaluating models only on naturally noisy data may accurately reflect their real-world performance, it may not inform about the specific type of perturbations the models are weak against, because noise sources present in these datasets are not controlled or even fully known. For example, the crowdsourced recordings in Common Voice [24] contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers. If a model struggles with Common Voice, identifying the specific types of noise it is sensitive to is not a straightforward task.
By evaluating models on a diverse set of known perturbations and noise sources, including adversarial attacks, SRB complements and extends benchmarks, like [23], that measure average case real-world performance and, thus, allows the robustness of ASR models to be comprehensively evaluated.
3 Speech Robust Bench
Figure 1: An illustration of the processes involved in using our benchmark to evaluate the robustness of ASR models.
At a high level, Speech Robust Bench (SRB) consists of four components: (1) Clean datasets, and a curated selection of inherently noisy datasets containing limited and known sources of noises and variations that are difficult to simulate; (2) Large bank of perturbations applicable on audio data; (3) Transcription extraction from ASR models and analysis; (4) Set of metrics for quantifying ASR model robustness, and facilitating comparisons among models. The basic workflow of SRB is the following (see Fig. 1): first, recordings from the clean dataset(s) (§ 3.1) are perturbed by a large bank of perturbations (§ 3.2), resulting in several perturbed versions of each recording. Next, these perturbed recordings, the original clean recordings, and the recordings with inherent noise are transcribed using the target ASR model. The predicted and reference transcripts are compared, and the accuracy and robustness of each model under each type of noise are captured with various metrics (§ 3.3). To account for the relative differences in the level of distortion introduced by different perturbations, we estimate speech quality scores using appropriate models and use them to normalize the metrics.
3.1 Dataset
We use clean speech from Librispeech [26] and TEDLIUM [27]. Librispeech contains professional recordings of audio books, and thus, has properly enunciated and cleanly recorded speech. Meanwhile, TEDLIUM contains professional recordings of TED talks and provides a degree of lexical and phonetic diversity that may be lacking in Librispeech. We chose these datasets due to their popularity in the ASR literature, however, any dataset containing utterances and reference transcripts would also be compatible with SRB. In addition to clean speech, SRB also uses inherently noisy speech containing specific types of perturbations that are hard to simulate, but are likely to be encountered by ASR models in the real-world. Specifically, SRB uses accented speech from Common Voice [24], and inter-personal conversations and far-field recordings from CHiME-6 [7] and AMI [8].
3.2 Perturbations
The taxonomy of the perturbations used in our benchmark is presented in Fig. 2. The perturbations are of two broad types: (1) non-adversarial and (2) adversarial. The non-adversarial perturbations fall under one of four categories: (1.1) Inter-personal communication gathered in social settings using near-field and far-field microphones. This setting is representative of the environments in which voice-enabled smart devices are expected to operate. The data for this category are drawn from the ChiME [7] and AMI [8] corpora. (1.2) Environmental effects, such as white noise, environmental noise, spatial acoustics (room impulse response and echo) (1.3) Digital augmentations produced by standard audio processing operations, and special effects. The audio processing operations included amplitude gain, resampling, lowpass, and highpass filtering. The special effects included bass gain, treble gain, tempo increase, tempo decrease, speed increase, speed decrease, pitch increase, pitch decrease, chorus, tremolo, and phaser. All special effects are applied using SoX.  [^1] https://sourceforge.net/projects/sox/ (1.4) Speech Variations are represented by speech from non-American speakers, and speech generated by a TTS model [59]. Since users from a variety of ethnic backgrounds may interact with ASR models, they must be robust to a variety of accents. With the rise of generative AI, robustness to synthetic speech will also be important.
Figure 2: Taxonomy of perturbations currently used in SRB.
The adversarial perturbations we use are untargeted adversarial attacks whose objective is to prevent the correct transcription of input speech. We use two types of adversarial attacks: (2.1) utterance-specific and (2.2) utterance-agnostic attacks. The utterance-specific attack optimizes a perturbation for a given speech recording, such that a given model maximally mistranscribes. The utterance-agnostic attack works similarly but the perturbation is jointly optimized over many speech recordings. Thus, this perturbation can affect several speech recordings, even those that were not used during the optimization. This represents a more realistic scenario where a malicious actor tries to mount a denial-of-service style attack against an ASR model by introducing utterance-agnostic perturbation at some point in the transcription pipeline.
Most perturbations, except for those categorized as informal speech, or speech variations, are applied at 4 levels of severity, which are modulated by adjusting the perturbation parameters. Details about the perturbations and their parameters are presented in Appendix A.
3.3 Metrics
SRB measures the utility of the model with the widely used Word Error Rate. WER is computed as the word-level edit distance between the reference and the predicted transcripts, normalized by the length of the reference (see Appendix B for formal definitions). To measure the robustness of the model to input perturbations, we use WER Degradation (WERD), computed as $WER(\mathcal{X}_{p})-WER(\mathcal{X})$, where $\mathcal{X}$ contains clean audio, while the audio in $\mathcal{X}_{p}$ is perturbed by perturbation $p$. If $p$ is synthetically applied, both $\mathcal{X}$ and $\mathcal{X}_{p}$ contain the same utterances, otherwise $\mathcal{X}$ contains clean audios from LibriSpeech and TEDLIUM.
When aggregating metrics (WER/WERD) over multiple perturbations, we follow the practice of [1] and divide the metric by a measure of difficulty, i.e., by the (estimated) speech quality degradation. This adds weight to errors on “easy” corruptions (less quality degradation) and under weighs errors on “harder” corruptions (more quality degradation) when computing averages. We refer to the difficulty normalized versions of WER/WERD as Normalized WER/WERD (NWER/NWERD). We estimate speech quality using DNSMOS [60] and PESQ[61], which are models of human judgments of speech quality and predict Mean Opinion Scores (MOS) [62]. PESQ uses various signal processing methods to predict MOS, while DNSMOS uses DNNs to do the same. To compute speech quality degradation we compute PESQ and DNSMOS for each clean and noisy recording multiplied by -1 (lower values indicate less degradation). Since we are only interested in the relative degradation between perturbations, we normalize the scores to have mean 50 and standard deviation 25.
4 Evaluation
We evaluate several recent and popular ASR DNNs (§4.1) using SRB and analyze the results to uncover fine-grained differences in their robustness to non-adversarial and adversarial input perturbations. We further extend our analysis by measuring ASR model robustness for various sub-groups, namely English speech and non-English (Spanish) speech, and male and female speakers. Prior works [63, 64] observe that there is a disparity in transcription accuracy between subgroups. Our analysis augments these observations by showing that inter-group disparities in robustness may also exist, thus demonstrating the utility of SRB in the broader field of trustworthy AI.
4.1 Models
For English speech, we evaluate Whisper[19] variants: large-v2 (wsp-lg), medium (wsp-md), small (wsp-sm), base (wsp-bs), and tiny (wsp-tn, wsp-tn.en), Wav2Vec-2.0 [38] variants: base (w2v2-bs), large (w2v2-lg), self-trained large [65] (w2v2-lg-slf), and Robust Wav2Vec [22], HuBERT [39] variants: large (hubt-lg) and XL (hubt-xl), Nvidia Canary [25] (cnry-1b), MMS [28] (mms-1b), Speech-T5 [66] (spch-t5) and DeepSpeech [67] (ds). For Spanish speech, we evaluate on the mono-lingual Wav2Vec base Spanish [68] (w2v2-bs-es), Wav2Vec XLSR Spanish [69] (w2v2-lg-es), wsp-lg, wsp-bs, wsp-tn, and mms-1b. We used the Huggingface implementations for all the models except cnry-1b (https://huggingface.co/nvidia/canary-1b) and ds (https://github.com/SeanNaren/deepspeech.pytorch). More details about the models are in Table 3 in the Appendix.
Figure 3: The accuracy and robustness of English (top) and Spanish (bottom) ASR models on clean and perturbed data. Accuracy is measured by WER of the models on clean speech (a & d). Robustness is measured by the NWERD on non-adversarially perturbed speech (b & e) and WERD on adversarially perturbed speech (c & f). The markers indicate the dataset in (a & c), and the perturbation category in (b & e). The x axes are in ascending order of the values on the y axes.
4.2 Accuracy on Clean Data
Figures 3(a) and 3(d) show the WER of the models on clean English and Spanish, respectively. We see that cnry-1b is the most accurate model on both languages, followed by w2v2-lg-slf on English and wsp-lg on Spanish.
4.3 Robustness of English ASR Models to Non-Adversarial Perturbations
Fig. 3(b) presents the NWERD for the English ASR models under non-adversarial noises and perturbations. We observe that wsp-lg emerges as the most robust model, followed by wsp-md and cnry-1b. Despite having the lowest overall WERD, wsp-lg is outperformed by wsp-md and cnry-1b on white noise, environmental noise, and special effects (see Appendix D for more detailed results). The massively multilingual mms-1b has the highest NWERD in social scenarios and far-field recordings category, indicating that it is unsuited for applications like voice assistants and smart home set-ups. We also notice that Wav2Vec-2.0 based models also struggle with the “speech variations” indicating that they are not suitable when users have diverse accents. Surprisingly, spch-t5 is the least robust model, and performs even worse than the much smaller and older ds. Similarly, wsp-md outperforms the newer and larger cnry-1b on almost all corruption categories.
Figure 4: WERD of cnry-1b, wsp-md and wsp-sm on perturbations in the spatial acoustics and special effects categories.
Upon closer examination, we see that cnry-1b has higher NWERD than its peers (wsp-md and wsp-sm) for two categories of perturbations: spatial acoustics and special effects. Upon looking at the WERD on each perturbation within these categories (Fig. 4) we find that cnry-1b is more sensitive than its peers to echo, real room impulse responses, and speed and pitch modifications. This analysis demonstrates that SRB can evaluate the robustness of ASR models at multiple ganularities and can pinpoint the weaknesses of a given model. Detailed results can be found in Figs. 7 and 9 in the appendix.
Takeaway: cnry-1b is the most accurate, but not most robust; Wav2Vec variants struggle against speech variations and social settings, thus, may not be suitable when users have diverse accents.
4.4 Robustness of English ASR Models to Adversarial Perturbations
Fig. 3(c) shows the WERD of the models on utterance specific and utterance agnostic adversarial attacks. We observe that hubt-lg and hubt-xl emerge as the most robust models to utterance-specific adversarial attacks, followed by other models based on Wav2Vec-2.0 large, Whisper, and Canary. Interestingly, Wav2Vec models that exhibited mediocre robustness to non-adversarial perturbations are robust to utterance-specific attacks, while Whisper models and Canary, which were the most robust on non-adversarial perturbations, struggle on utterance specific attacks. We also note that most Wav2Vec models are considerably more robust to attacks against TEDLIUM than against LibriSpeech, and the opposite is true for whisper and Canary models. Under utterance-agnostic attacks, the best-performing models are mms-1b, wsp-bs, and wsp-sm. It is interesting to note that the smaller variants of Whisper limit the generalizability across utterances of the adversarial perturbations to a greater extent than their larger counterparts.
Takeaway: Models that are most robust to non-adversarial perturbations, are mediocre against adversarial perturbations; cnry-1b is highly vulnerable to utterance specific attacks.
4.5 Robustness vs. Model Attributes
To determine if the prevailing practice of training DNNs with more parameters on larger datasets is yielding improvements in robustness, we use robust linear regression to fit a line to WERD vs. number of model parameters/size of the training data for the candidate models in Figs. 5(a) and 5(b), respectively. Increasing model size is correlated with improved robustness (lower WERD).To further isolate the impact of the model size we plot the NWERD of models from the same family in Fig. 5(c), which have similar architectures and training datasets. We note that larger models are more robust in the Whisper and Wav2Vec-2.0 families, but, surprisingly, not in the HuBert family. In contrast, increasing training data appears to have only a minor influence on robustness (Fig. 5(b)).
Takeaway: Larger models tend to be more robust, while smaller models, even if they are trained on large datasets, are less robust. This runs somewhat counter to the prevailing wisdom [19, 22].
Figure 5: (a & b) WERD for all models with robust regression fitted line on non-adversarial (blue) and adversarial (orange) perturbations, plotted against (a) number of parameters, and (b) hours of training data. (c) NWERD lineplot with non-adversarial and adversarial perturbations, three families of models (color indicated). Models per family share similar architectures and training datasets. (d) Robustness of multi-lingual English (solid) and Spanish (hashed) models.
4.6 Robustness for Population Sub-Groups
In the preceding analysis, we considered robustness aggregated over the entire population (i.e., dataset). However, populations are generally not homogeneous, and, thus, the robustness of the model may differ on various population sub-groups. Prior works have commonly analyzed sub-group fairness of ASR models by comparing the overall WER for each sub-group on a benchmark dataset [70]. It is possible that models that are fair on average, may not be fair under certain conditions. In the following, we use SRB to uncover and analyze the disparities in the models’ robustness across four sub-groups: English and Spanish speech, and male and female speakers. We find that disparities indeed exist, with multi-lingual models generally being more robust for English than Spanish (Fig. 5(d)), and most models being less robust for females than males.
4.6.1 Robustness on Non-English (Spanish) Speech
The clean accuracy of the Spanish ASR models is presented in Fig. 3(d), and the NWERD under non-adversarial perturbations for each model is shown in Fig. 3(e). We observe, that the two most accurate models, cnry-1b and wsp-lg, are also the two most robust models, but mms-1b, which had fairly high WER, is the third most robust model in terms of NWER. We also notice a general trend that all the models, except cnry-1b, struggle against “spatial acoustics” category, which is rather worrying since the influence of spatial acoustics is ubiquitous in the real-world.
Considering adversarial attacks (Fig. 3(f)), we observe that Wav2Vec models exhibit better robustness than Whisper and Canary under both utterance-agnostic and utterance-specific perturbations. This is similar to the trends observed in English speech (Fig. 3(c)), and indicates that perhaps learned front-ends may be more adversarial robust than the spectrograms used by Whisper and Canary. Detailed results can be found in Fig. 8 in the appendix.
We also compare the robustness exhibited by multi-lingual models, wsp-lg, wsp-bs, cnry-1b and mms-1b on English and Spanish. The WERD of these models on both languages is presented in Fig. 5(d). We observe that Whisper models achieve lower WERD on English speech than on Spanish on almost all perturbation categories, while cnry-1b and mms-1b achieve similar WERD on some categories. We also note that the difference in WERD on some common perturbation categories, like environmental noise, and spatial acoustics, is much greater for wsp-lg than for cnry-1b.
Takeaway: Multilingual models are more robust on English than Spanish; cnry-1b and wsp-lg most robust on both languages; adversarial robustness results follow the same trend as English.
4.6.2 Robustness and Gender
To measure the disparity in prediction accuracy across genders (males/females), we compute the log of the ratio of the WERs of the ASR model on female and male speakers. We call this measure the Log WER Ratio (LWERR). A positive value of LWERR indicates that the model is biased against females and a negative value indicates that the model is biased against males.
The LWERR for each dataset is shown in Fig. 6. We note that, on average, the models are biased against females on LibriSpeech and Spanish Multilingual Librispeech (MLS-ES), and against males on TEDLIUM, Common Voice and AMI. The bias is most prominent in MLS-ES, where cnry-1b seems to be yielding the highest disparities among genders. We also note that adversarial perturbations cause the WER of wsp-lg to increase significantly more for females than males in LibriSpeech. This is interesting because adversarial perturbations do not target a specific part of the spectrum and thus should not impact one gender more than the other.
Takeaway: Models are more robust for males on some datasets, and females on other datasets suggesting that used data require further examination; adversarial attacks increase WER of Whisper variants for females more than males; multilingual models, particularly cnry-1b, are more biased against females when transcribing Spanish.
Figure 6: Log WER Ratio for various TEDLIUM, LibriSpeech, Common Voice (CV) and AMI, and Spanish Multilingual Librispeech (MLS (ES)). WERs are averaged across severity levels and individual augmentations within each category before computing the Log WER Ratio.
5 Limitations
Despite the comprehensive design of SRB, there are limitations to consider. Firstly, while SRB includes a diverse set of datasets, distortions and adversarial attacks, it may not encompass all possible real-world scenarios. Additionally, the benchmark has been tested for now on English and Spanish languages, which may not capture the robustness challenges faced by ASR models in other languages and dialects. While these limitations might potentially affect the generalizability of the results obtained with SRB, its extensibility facilitates the incorporation of novel approaches for quantifying ASR robustness.
6 Conclusion
We propose SRB, a comprehensive benchmark designed to standardize the robustness evaluations of Automatic Speech Recognition (ASR) models. To the best of our knowledge, SRB is the first ASR robustness benchmark that contains long-form speech, and expands the set of distortions to include special effects, speaker attribute modification and adversarial attacks. We demonstrate the utility of SRB in evaluating the robustness of ASR models via several concrete examples, as well as its potential to facilitate evaluations of other aspects of trustworthy AI, like fairness. We believe that SRB will enable rigorous robustness evaluations of ASR models in a highly standardized manner, allowing easy comparisons between existing and new approaches. To further facilitate robustness evaluations for researchers and model developers, we release transformed test sets in English and Spanish. We anticipate that this will make robustness evaluations more prevalent and encourage model developers to consider robustness as a key metric for improvement.
Appendix A Perturbation Generation/Application Procedure
Below, we provide further details the perturbations that make up Speech Robust Bench. Table 1 shows the parameters for each perturbation and Table 2 shows the normalized DNSMOS and PESQ scores for each perturbation.
Category Perturbation Sev 1 Sev 2 Sev 3 Sev 4
Gaussian Noise 30 dB 20 dB 10 dB 0 dB
Environment Environmental Noise 30 dB 20 dB 10 dB 0 dB
Music 30 dB 20 dB 10 dB 0 dB
Crosstalk 30 dB 20 dB 10 dB 0 dB
Spatial Acoustics RIR 0.27s 0.58s 0.99s 1.33s
Real RIR 9.1 7.1 4.1 1.8
Echo (delay) 125 ms 250 ms 500 ms 1000 ms
Special Effects Bass (gain) 20 30 40 50
Treble (gain) 10 23 36 50
Phaser (decay) 0.3 s 0.5 s 0.7 s 0.9 s
tempo-up 1.25x 1.5x 1.75x 2x
tempo-down 0.875x 0.75 0.625x 0.5x
Speed-up 1.25x 1.5x 1.75x 2x
Slow-down 0.875x 0.75 0.625x 0.5x
Pitch Step-up 0.25 oct 0.5 oct 0.75 oct 1 oct
Pitch Step-down 0.25 oct 0.5 oct 0.75 oct 1 oct
Chorus (delay) 30 50 70 90
tremolo (depth) 50 66 83 100
Audio Processing Resampling 0.75x 0.5x 0.25x 0.125x
Gain (factor) 10x 20x 30x 40x
Low-pass filter 4 kHz 2833 kHz 1666 kHz 500 kHz
High-pass filter 500 kHz 1333 kHz 2166 kHz 3000 kHz
Adversarial PGD Attack 40 dB 30 dB 20 dB 10dB
Utterance Agnostic Attack 40 dB 30 dB 20 dB 10dB
Table 1: The parameters defining the various severity levels of the perturbations used in the proposed benchmark.
Metric $\rightarrow$ AVG normalized DNSMOS normalized PESQ
Perturbation / Severity 1 2 3 4 1 2 3 4 1 2 3 4
accent 31.6 - - - 31.6 - - - - - - -
bass 18.3 23.0 35.0 55.2 22.5 27.4 38.0 56.7 11.6 14.5 30.5 54.6
chorus 39.1 48.2 54.4 55.9 29.8 39.7 47.1 48.2 63.7 71.1 74.2 75.7
crosstalk 22.3 38.2 52.3 59.1 23.3 31.6 36.7 39.4 22.4 46.2 70.0 81.4
echo 54.1 53.4 52.8 50.6 39.0 36.3 36.0 34.9 71.2 72.6 71.7 67.9
env noise 50.5 61.5 76.0 88.5 51.8 59.2 72.7 91.4 39.7 58.7 76.6 84.1
env noise esc50 26.1 40.9 57.5 72.8 37.7 43.7 53.3 71.3 20.3 43.2 65.9 79.7
env noise musan 24.3 42.3 62.1 75.4 24.8 37.2 55.3 71.2 25.0 48.7 70.3 80.9
env noise wham 22.4 45.4 73.2 92.2 22.3 39.7 70.5 99.5 23.3 52.0 76.5 85.2
gain 50.0 68.9 76.6 80.7 45.1 65.9 76.4 82.6 61.3 75.9 79.8 81.6
gnoise 52.4 75.6 90.5 100.9 70.3 87.5 101.6 117.3 42.2 69.1 82.8 86.1
highpass 40.2 55.4 67.5 77.5 34.3 43.4 64.9 81.0 46.2 68.3 71.5 74.8
itw ff 100.1 - - - 100.1 - - - - - - -
itw ff ami 83.9 - - - 83.9 - - - - - - -
itw nf 78.1 - - - 78.1 - - - - - - -
itw nf ami 35.8 - - - 35.8 - - - - - - -
lowpass 33.1 37.1 50.8 78.0 47.3 46.7 62.5 97.9 20.5 29.3 40.5 58.4
music 22.3 43.1 65.8 78.9 25.5 41.4 62.2 76.7 20.2 45.4 70.0 81.7
phaser 15.0 32.3 59.8 79.5 21.4 34.7 58.1 76.7 10.9 32.0 63.5 83.4
pitch down 60.9 67.3 53.4 83.3 38.0 49.2 65.5 80.8 85.7 86.3 55.5 86.6
pitch up 58.0 61.2 64.3 65.1 32.2 38.1 44.1 45.9 85.7 86.2 86.3 86.3
real rir 38.7 53.8 68.9 84.2 34.2 44.8 59.9 87.1 43.2 62.7 77.9 81.3
resample 14.4 27.3 49.0 63.3 23.8 42.1 61.5 75.5 7.0 18.3 38.2 52.5
rir 50.3 63.1 68.3 68.0 40.4 56.3 64.3 64.7 64.9 74.5 78.0 78.0
slowdown 50.7 57.0 64.3 72.7 17.6 29.7 43.6 66.7 85.7 85.9 85.8 86.2
speedup 51.5 58.7 66.3 72.9 21.7 36.1 50.9 63.7 83.6 83.4 83.1 82.8
tempo down 48.7 51.8 54.5 50.1 18.4 21.6 26.7 35.1 81.3 84.5 85.1 85.5
tempo up 50.2 57.1 63.2 69.7 24.3 34.6 46.1 58.3 78.9 82.1 82.3 82.3
treble 11.6 21.6 40.5 62.4 19.4 29.7 43.2 60.6 2.4 12.4 42.3 72.4
tremolo 16.8 29.0 59.0 98.7 23.1 37.2 72.1 111.4 9.9 17.8 37.1 75.5
universal adv 10.3 - - - 11.7 - - - 8.8 - - -
voice conversion vctk 49.6 - - - 15.7 - - - 83.4 - - -
Table 2: Normalized DNSMOS and PESQ score for each perturbation.
Gaussian Noise:
A noise vector of the same length as the audio signal is sample from a standard normal distribution, scaled such that its magnitude corresponds to a specific SNR, and then added to the audio signal. We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
Environmental Noise:
We use the recordings of environmental noises from the test/eval subsets of ESC-50  [71], MS-SNSD [60], MUSAN [14] and WHAM [9]. We create a separate perturbed version of the clean data using each of these noise datasets. To do so, for each test utterance we sample a random environmental noise and add it to the audio signal at the specified SNR. We clip the noise if it is longer than the speech, and repeat it if it is shorter than the speech. We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
Room Impulse Response:
The simulated and real RIRs from [16] are applied to clean recordings. As a measure of intensity, RT60 is estimated for the simulated RIRs using Sabine’s formula with the room dimensions and absorption coefficient provided in the dataset. For the real RIRs, we compute the SRMR [72] using the implementation from https://github.com/aliutkus/speechmetrics/tree/master. The severity is defined in increasing RT60s for the synthetic RIRs, and decreasing SRMR for the real RIRs. Table 1 shows the average RT60/SRMS in each severity level. During evaluation, a random RIR having the given severity level is sampled for each test recording.
Resampling, Speed, Pitch, and Gain Perturbations:
The resampling speed, pitch, and gain perturbations were applied using the Resample Speed, PitchShift and Vol transforms from torchaudio.
Other special effects:
These effects are applied via SoX filters of the same name. We used torchaudio.sox_effects.apply_effects_tensor to apply these filters to the audio. The args for each filter are as follows:
  • echo 0.8 0.9 <delay> 0.3
  • phaser 0.6 0.8 3 <decay> 2 "-t"
  • Tempo <factor> 30
  • sinc <lo-freq>
  • sinc 0-<hi-freq>
  • tremolo 20 <depth>
  • treble <gain>
  • bass <gain>
  • chorus 0.9 0.9 <delay> 0.4 0.25 2 -t {<delay>+10} 0.3 0.4 2 -s
Voice Conversion
We use use YourTTS [59] from Coqui.ai [^2] https://github.com/coqui-ai/TTS to synthesize audio from textual transcripts in a given speaker’s style. The transcripts from the test clean subset of LibriSpeech are used. The target speakers are drawn from the VCTK corpus [73], which contains accented speech from 12 accents. For each transcript a random speaker is chosen to synthesize the audio.
Untargeted Adversarial Attack:
The adversarial perturbations are computed using an untargeted PGD adversarial attack. The attack is computed as follows. First, the maximum possible L2 norm of the noise is determined by solving the equation for SNR for the norm of the noise.
[…]
where $\delta$ is the noise, $x$ is the audio signal and SNR is the upper bound on the SNR in the final signal. Then, we follow the approach of [43] and optimize $\delta$ using Projected Gradient Descent (PGD) to maximize the divergence between the true and predicted transcriptions.
A.1 Crosstalk and Music
We use crosstalk and music audios from MUSAN [14]. We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
A.2 Accents
We select a subset of audios from the test set of Common Voice 17. The selected audios satisfied the following criteria: (1) the speaker’s accent must be present in the metadata, (2) the accent must not be American, (3) should be clean. The last criterion is satisfied if the DNSMOS [10] score of the recording is at least 3.4. The resulting subset contains 640 recordings. The most popular accent in this set is South Asian (India, Pakistan, Sri Lanka) (2̃5%), followed by British English (2̃5%).
A.3 Inter-Personal Communications
We CHiME-6 [7] and AMI [8] to obtain recordings of people in social scenarios (dinner party and meetings). In both these datasets, the speakers are recorded through lapel microphone and a room microphone resulting in near and far field recordings. We use both types of recordings and show separate results for them. We remove recordings that contain less than three words since they are often fillers.
Untargeted Utterance Agnostic Adversarial Attack:
We use the method of [52] to compute utterance agnostic adversarial perturbations. Concretely, we compute a single perturbation vector that can cause the accuracy of the model on several (ideally all) inputs to degrade severely. The main difference between the universal attack and the PGD attack is that the latter computes a perturbation vector for each input, whereas the former computes a single perturbation that is expected to successfully attack any input to the model. Once we compute the perturbation we add it to the audio signal at the specified SNR.
We use the robust_speech package [74] to generate adversarial perturbations.
[…]
Appendix B Additional definitions
Word Error Rate:
As noted in the main text, we use word error rate (WER) as a basic measure for quantifying performance of the models. Following the common practice from ASR literature, the WER is computed as the word-level edit distance between the reference and the predicted transcripts, normalized by the length of the reference. The edit distance is computed as the total number of word substitutions, deletions, and additions required to transform the reference transcript into the predicted transcript. We remove all punctuation from both the reference and predicted transcripts, and convert to lower case before computing WER.
When WER is computed over multiple pairs of predicted and reference transcripts, it is common practice to sum the number of substitutions, deletions, and additions for all the pairs, and divide by the sum of the lengths of the reference transcripts. Formally, this can be written as:
$WER^{M}(\mathcal{X},\mathcal{R}):=100\frac{\sum_{x\in\mathcal{X},r\in\mathcal{R}}ED(M(x),r)}{\sum_{r\in\mathcal{R}}|r|}\%,$ (4)
where $ED$ computes the edit distance.
For quantifying differences between (binary) genders, we measure the disparity in prediction accuracy across males and females by the Log WER Ratio (LWERR). Formally,
$LWERR:=\log_{2}\frac{WER^{M}(\mathcal{X}_{f},\mathcal{R}_{f})}{WER^{M}(\mathcal{X}_{m},\mathcal{R}_{m})},$ (5)
where the $(\mathcal{X}_{f},\mathcal{R}_{f})$ and $(\mathcal{X}_{m},\mathcal{R}_{m})$ represent the subsets of utterances by females and males respectively.
Fairness through robustness:
In this work, we use SRB to conduct a fairness assessment based on robustness disparities across population subgroups (English vs. Spanish speech; male vs. female speakers). Most of the state of the art quantifies fairness in terms of predictive performance disparities. This occurs in the domain of fairness in ASR systems [63, 64, 70] and similar prevalence is found in other domains [75, 76]. However, previous work in the domain of ASR also considered robustness disparities as an alternative fairness notion [77].
Moreover, we argue that considering the dimension of robustness could give better sense of the expected disparities that could be observed when deployed in the wild, in the presence of diverse noisy conditions.
Appendix C Models
Table 3 provides a summary of the models used in our evaluations. The model names correspond to the names of their pretrained checkpoints in the Huggingface library (https://huggingface.co/models). The abbreviations of these names are in the parentheses after them. Some of the unilingual models are pre-trained on multilingual data but are fine-tuned on only one language and thus can not transcribe any other language. Multilingual models have been pre-trained and fine-tuned on multiple languages so the same DNN can transcribe several languages. The WER of multilingual models is presented as English/Spanish.
# Params (M) Data (hrs) WER
Language Model model (abrv)
EN canary-1b [25] cnry-1b 1,000.0 85,000 6.0
deepspeech[67] ds 86.0 960 26.5
hubert-large-ls960-ft[39] hubt-lg 317.0 60,000 8.4
hubert-xlarge-ls960-ft[39] hubt-xl 964.0 60,000 8.4
mms-1b-fl102 [78] $m$s-1b 964.6 55,000 22.8
speecht5 asr[66] spch-t5 154.6 960 22.1
wav2vec2-base-960h [38] w2v2-bs 95.0 960 11.3
wav2vec2-large-960h [38] w2v2-lg 317.0 960 9.7
wav2vec2-large-960h-lv60-self [65] w2v2-lg-slf 317.0 60,000 7.7
wav2vec2-large-robust-ft-libri-960h [58] w2v2-lg-rob 317.0 63,000 8.9
whisper-base [19] wsp-bs 74.0 680,000 9.6
whisper-base.en [19] wsp-bs.en 74.0 563,000 5.1
whisper-large-v2 [19] wsp-lg 1,550.0 680,000 8.0
whisper-medium [19] wsp-md 769.0 680,000 7.9
whisper-medium.en [19] wsp-md.en 769.0 563,000 4.1
whisper-small [19] wsp-sm 244.0 680,000 8.3
whisper-small.en [19] wsp-sm.en 244.0 563,000 4.0
whisper-tiny [19] wsp-tn 39.0 680,000 11.3
whisper-tiny.en [19] wsp-tn.en 39.0 563,000 10.1
ES canary-1b [25] cnry-1b 1,000.0 85,000 3.2
mms-1b-fl102 [78] $m$s-1b 964.6 55,000 15.7
wav2vec2-base-10k-voxpopuli-ft-es [68] w2v2-bs-es 94.4 10,116 25.7
wav2vec2-large-xlsr-53-spanish [69] w2v2-lg-es 315.4 54,350 6.8
whisper-base [19] wsp-bs 74.0 680,000 14.8
whisper-large-v2 [19] wsp-lg 1,550.0 680,000 5.8
whisper-tiny [19] wsp-tn 39.0 680,000 23.3
Table 3: Models used in our evaluations.
Appendix D Fine Grained Analyses
The following figures present fine-grained analyses of robustness. These figures may be referenced by the main text but were not included in the main body in the interest of space. Figures 7, and 8 present the breakdown by perturbation of the robustness of models on English and Spanish, respectively. Figure 9 presents a breakdown of robustness by severity.
Figure 7: NWERD of English models on different augmentations, averaged over all severities.
Figure 8: NWERD of Spanish models on different augmentations, averaged over all severities.
Figure 9: NWERD on English data as the severity of the augmentation is increased.
Figure 10: Log WER Ratio between male and female speakers from Librispeech (English) (a) and Spanish Multilingual Librispeech (b).
Appendix E Compute Resources
The experiments were performed on the Bridges-2 cluster at the Pittsburgh Supercomputing Center. This cluster contains 200 32G and 16G Nvidia V-100, which were used for these experiments.
Appendix F Dataset Licenses
The licenses of each of the considered datasets are described in Table 4.
Dataset License
LibriSpeech $C$-BY-4.0
Multilingual Librispeech CC BY 4.0
TEDLIUM CC-BY-NC-ND 3.0
AMI $C$-BY-4.0
Common Voice CC0-1.0
CHiME CC BY-SA 4.0
Table 4: Licenses of each of the considered datasets in SRB

### v3

@@
 We propose Speech Robust Bench (SRB), a comprehensive benchmark for evaluating the robustness of ASR models to diverse corruptions.
-SRB is composed of 114 input perturbations which simulate an heterogeneous range of corruptions that ASR models may encounter when deployed in the wild.
+SRB is composed of 114 challenging speech recognition scenarios which largely cover the range of corruptions that ASR models may encounter when deployed in the wild.
 We use SRB to evaluate the robustness of several state-of-the-art ASR models and observe that model size and certain modeling choices such as the use of discrete representations, or self-training appear to be conducive to robustness.
@@
 1 Introduction
-As novel ML models continue to be developed and deployed at an ever-increasing rate, it has become crucial to ensure that their robustness to corruptions arising from a myriad of sources, including the environment, sensing apparatus, and even malicious actors is well understood.
-To this end, prior works have developed comprehensive robustness benchmarks, particularly for vision [1, 2, 3, 4] and natural language processing models [5, 6], that evaluate several aspects of a model’s performance under a variety of input perturbations.
-These benchmarks have proven to be invaluable to the advancement of research into more robust models because (1) they unify robustness evaluations, thus enabling meaningful comparisons across models and allowing progress to be accurately tracked, and (2) they make it easier for researchers to comprehensively evaluate the robustness of their models by aggregating a diverse and representative set of perturbations in a single benchmark.
-While several robustness benchmark datasets exist for Automatic Speech Recognition (ASR) models [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], none of the currently existing ones is in any sense comprehensive: each benchmark measures the model’s robustness w.r.t.
-to one or a few specific types of perturbations, thereby putting the onus on model developers to find and collect all the relevant benchmarks to evaluate their model comprehensively.
-This has often resulted in model developers evaluating their models on disparate benchmarks  [19, 20, 21, 22], which makes it hard to reliably compare performance and robustness across models.
-Recently, Huggingface Open ASR Leaderboard [23] has sought to unify ASR model evaluations by developing a benchmark consisting of several real-world speech datasets.
-Although evaluating models on exclusively natural data may accurately reflect their real-world performance, it is generally not informative about the specific types of perturbations the models are weak or strong against, because the noise sources present in these datasets are not controlled or even fully known.
-For example, the crowdsourced recordings in Common Voice [24] contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers.
-Additionally, crucial digital perturbations like special effects, computer-generated speech, and adversarial examples, to which humans are usually invariant, are largely overlooked by existing benchmarks.
+As novel ML models continue to be developed and deployed at an ever-increasing rate, it has become crucial to ensure their robustness to challenging real-world scenarios, where corruptions arising from a myriad of sources, including the environment, sensing apparatus, and even malicious actors are present.
+To this end, prior works have developed comprehensive robustness benchmarks, particularly for vision (Hendrycks & Dietterich, 2019; Hendrycks et al., 2021a; b; Croce et al., 2020) and natural language processing models (Wang et al., 2021a; 2022b), that evaluate a model’s performance under a variety of challenging scenarios.
+These benchmarks have proven to be invaluable to the advancement of research into more robust models because (1) they unify robustness evaluations, thus enabling meaningful comparisons across models and allowing progress to be accurately tracked, and (2) they make it easier for researchers to comprehensively evaluate the robustness of their models by aggregating a diverse and representative set of scenarios, and methods of simulating them, in a single benchmark.
+While several robustness benchmark datasets exist for Automatic Speech Recognition (ASR) models (Barker et al., 2017; Kraaij et al., 2005; Wichern et al., 2019; Reddy et al., 2020; Cosentino et al., 2020; Hershey et al., 2016; Chen et al., 2020; Snyder et al., 2015; Kinoshita et al., 2013; Ko et al., 2017; Nakamura et al., 2000; Jeub et al., 2009), none of the currently existing ones are in any sense comprehensive, because each benchmark measures the model’s robustness w.r.t.
+to one or a few specific types of corruptions or scenarios, which puts the onus on model developers to find and collect all the relevant benchmarks to evaluate their model comprehensively.
+This has often resulted in model developers evaluating their models on disparate benchmarks  (Radford et al., 2023; Wen et al., 2016; Chen et al., 2022; Likhomanenko et al., 2020), which makes it hard to reliably compare performance and robustness across models.
+Recently, Huggingface Open ASR Leaderboard (Srivastav et al., 2023) has sought to unify ASR model evaluations by developing a benchmark consisting of several real-world speech datasets.
+Although evaluating models on exclusively natural data may accurately reflect average case real-world performance, it is generally not informative about the specific types of corruptions the models are weak against, because the noise sources present in these datasets are not controlled or even fully known.
+For example, the crowdsourced recordings in Common Voice (Ardila et al., 2019) contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers.
+Furthermore, digital perturbations like special effects, computer-generated speech, and adversarial examples, that may be prevalent in digital content are largely overlooked by existing benchmarks.
 In this paper, we propose Speech Robust Bench (SRB), a benchmark for comprehensively evaluating the robustness of ASR models to input perturbations and corruptions.
-SRB is designed to address the aforementioned major shortcomings of existing ASR robustness benchmarks, i.e., that (1) they are often specialized and thus are not individually comprehensive, (2) even taken together, they overlook important perturbation types.
-SRB addresses these shortcomings by curating a large comprehensive bank of 114 perturbations and inherently noisy data sources.
-The perturbation bank includes distortions and variations caused by the environment, recording equipment, special effects, computer-generated speech, and significantly, adversarial attacks, which are often overlooked by existing benchmarks.
-When selecting inherently noisy data sources, we take special care to ensure that the sources of variation in each data source are limited and known, so that the models’ specific weaknesses can be identified.
-These sources contain speech with non-native English accents and inter-personal conversations, giving rise to articulatory and lexical variability in speech.
+SRB is designed to address the aforementioned major shortcomings of existing ASR robustness benchmarks, i.e., that (1) they are often specialized and thus are not individually comprehensive, (2) even taken together, they overlook important challenging scenarios, like special effects and adversarial attacks, and (3) may not reveal the specific weaknesses of the models.
+SRB addresses these shortcomings by evaluating ASR models under a comprehensive set of challenging scenarios, using recordings that are either are recorded under specific scenarios, and thus are inherently “noisy”, or recordings that are digitally perturbed to simulate the various scenarios.
+SRB uses real recordings of accented speech and inter-personal conversations to evaluate robustness to articulatory and lexical variability.
+We take care to ensure that the recordings are clean and do not have any other corruption that may confound the results.
+To digitally simulate challenging scenarios, we curate a large comprehensive bank of 114 perturbations that represent common distortions arising from the environment, recording equipment, special effects, computer-generated speech, and adversarial attacks that are often overlooked by existing benchmarks.
 To highlight the need for and the benefits of doing systematic and comprehensive robustness assessment, we evaluate the robustness of several popular ASR models (§4.1) using SRB.
-We observe that Whisper [19] is the most robust on average among the models we tested, even outperforming the more recent Canary [25].
+We observe that Whisper (Radford et al., 2023) is the most robust on average among the models we tested, even outperforming the more recent Canary (NVIDIA, ).
 We conduct further analyses to disentangle the effects of model and training data size, revealing that larger models tend to be more robust than smaller models, even if the latter are trained on significantly more data.
@@
 This highlights the broad utility of this benchmark to the field of trustworthy AI.
-  • To facilitate out-of-the-box robustness evaluations for the research community, we have publicly released a large dataset (https://huggingface.co/datasets/mshah1/speech_robust_bench) containing perturbed versions of LibriSpeech [26], TEDLIUM [27] and Spanish Multilingual LibriSpeech [28] test sets, and accented speech from common voice and segemented near- and far-field audios from CHiME-6 [10] and AMI[8].
-  • We open source our code with clear documentation to allow for easy reproducibility and extensibility.(https://github.com/ahmedshah1494/speech_robust_bench)
+  • To facilitate out-of-the-box robustness evaluations for the research community, we have publicly released a large dataset (link omitted) containing perturbed versions of LibriSpeech (Panayotov et al., 2015), TEDLIUM (Hernandez et al., 2018) and Spanish Multilingual LibriSpeech (Pratap et al., 2020) test sets, and accented speech from common voice and segemented near- and far-field audios from CHiME-6 (Reddy et al., 2020) and AMI(Kraaij et al., 2005).
+  • We open source our code with clear documentation to allow for easy reproducibility and extensibility. (https://anonymous.4open.science/r/speech_robust_bench-4C77)
 2 Related Work
 2.1 Robust Automatic Speech Recognition
-Several techniques have been proposed for making Automatic Speech Recognition (ASR) models robust to input perturbations, such as noise and other signal corruptions [29].
+Several techniques have been proposed for making Automatic Speech Recognition (ASR) models robust to input perturbations, such as noise and other signal corruptions (Li et al., 2014).
 We can divide these techniques into two high-level categories: i) model-based and ii) feature-based.
 Model-based techniques modify the models to make them more robust.
-Examples of such approaches include adapting pre-trained models [30, 31], de-noising the audio before processing [32, 33], and training ASR models on noisy data [22].
-Since model-based strategies generally require access to noisy data [29], they are most effective if the sources of noise, and/or the exact environment in which the ASR model will be deployed in are known, and one can gather data to represent them.
-Feature-based approaches, on the other hand, involve developing handcrafted features, invariant to noise and corruptions in the signal [29].
-Several of these features are inspired by biological audition [34, 35, 36], while others use signal processing techniques [29].
-Generally, these methods are designed to extract the components of the audio signal salient for speech production and perception, while discarding irrelevant components [37].
+Examples of such approaches include adapting pre-trained models (Yu et al., 2009; Juang & Rahim, 1996), de-noising the audio before processing (Mohammadiha et al., 2013; Wilson et al., 2008), and training ASR models on noisy data (Likhomanenko et al., 2020).
+Since model-based strategies generally require access to noisy data (Li et al., 2014), they are most effective if the sources of noise, and/or the exact environment in which the ASR model will be deployed in are known, and one can gather data to represent those.
+Feature-based approaches, on the other hand, involve developing handcrafted features, invariant to noise and corruptions in the signal (Li et al., 2014).
+Several of these features are inspired by biological audition (Kim & Stern, 2016; Hermansky et al., 1991; Hermansky & Sharma, 1998), while others use signal processing techniques (Li et al., 2014).
+Generally, these methods are designed to extract the components of the audio signal salient for speech production and perception, while discarding irrelevant components (Stern & Morgan, 2012).
 Consequently, they do not require precise knowledge of the environment and noise distributions.
-Recently, however, handcrafted features have fallen out of favor, and have been replaced by features learned via end-to-end training of deep learning models on large amounts of data [38, 39, 22, 19].
+Recently, however, handcrafted features have fallen out of favor, and have been replaced by features learned via end-to-end training of deep learning models on large amounts of data (Baevski et al., 2020; Hsu et al., 2021a; Likhomanenko et al., 2020; Radford et al., 2023).
 Proponents of these techniques posit that models trained on larger datasets become more robust.
@@
 2.2 Adversarial Robustness
-Adversarial perturbations can change the response of a model when added to their inputs, but are either imperceptible to humans or perceptually and semantically irrelevant enough to be ignored by them [40, 41].
+Adversarial perturbations can change the response of a model when added to their inputs, but are either imperceptible to humans or perceptually and semantically irrelevant enough to be ignored by them (Szegedy et al., 2014; Goodfellow et al., 2014).
 Adversarially perturbed inputs are known as adversarial attacks.
-They can be targeted (aiming to change a prediction to a specific incorrect class), or un-targeted (aiming to change a prediction to any incorrect class [42]).
+They can be targeted (aiming to change a prediction to a specific incorrect class), or un-targeted (aiming to change a prediction to any incorrect class (Akhtar et al., 2021)).
 The design of adversarial attacks is determined by the level of knowledge the attacker is assumed to have about the target model.
-Attacks that assume full knowledge of the target model’s architecture and weights (white-box threat model) often use gradient-based optimization techniques [40, 41, 43, 44, 42].
-Attackers who do not have any knowledge of the target model’s architecture and only have query access to it (black-box threat model) typically use gradient-free optimization methods [45, 46, 47, 48, 49, 50].
-An intriguing property of adversarial perturbations is that they transfer between models [51], and inputs [42, 52], i.e.
+Attacks that assume full knowledge of the target model’s architecture and weights (white-box threat model) often use gradient-based optimization techniques (Szegedy et al., 2014; Goodfellow et al., 2014; Madry et al., 2018; Laidlaw et al., 2021; Akhtar et al., 2021).
+Attackers who do not have any knowledge of the target model’s architecture and only have query access to it (black-box threat model) typically use gradient-free optimization methods (Wang et al., 2022a; Andriushchenko et al., 2020; Wicker et al., 2018; Chen et al., 2017; Zhao et al., 2020; Vo et al., 2022).
+An intriguing property of adversarial perturbations is that they transfer between models (Papernot et al., 2016), and inputs (Akhtar et al., 2021; Neekhara et al., 2019), i.e.
 perturbations designed for one model/input may be effective against others as well.
-Our SRB includes two types of white box adversarial attacks; those that generate perturbations: 1) specific to each input [43], 2) that cause models to mis-transcribe multiple inputs [52].
+Our SRB includes two types of white box adversarial attacks; those that generate perturbations: 1) specific to each input (Madry et al., 2018), 2) that cause models to mis-transcribe multiple inputs (Neekhara et al., 2019).
 2.3 Robustness Benchmarks for Speech
-Several robustness benchmark datasets exist for Automatic Speech Recognition (ASR) models [7, 8, 9, 10, 11, 12, 9, 13, 14, 15, 16, 17, 18, 53, 54, 55].
-Unfortunately, these benchmarks do not comprehensively evaluate model robustness.
-While past works such as [54] did indeed propose benchmarks containing diverse perturbations, many of the recently proposed benchmarks measure the model’s robustness to one or a few specific types of perturbations.
-For example, some datasets [15, 17, 18, 16] focus exclusively on reverberant speech, while others focus only on multi-speaker scenarios [8, 7], environmental noise [14, 10, 9, 53], or accented speech [56, 57].
-In this situation, the onus is on model developers to collect all relevant benchmarks and comprehensively evaluate their model.
+While several robustness benchmark datasets exist for ASR models, they, unfortunately, suffer from three major shortcomings that make it difficult to perform robustness evaluations in a way that is comprehensive, sufficiently fine-grained, and comparable across models.
+Firstly, we find that existing benchmarks do not comprehensively evaluate robustness.
+While past works such as (Pearce & Picone, 2002) did indeed propose benchmarks containing diverse perturbations, they were limited to relatively simple data unsuitable for modern ASR models.
+However, many recently proposed benchmarks measure the robustness in one or a few specific types of scenarios.
+For example, some datasets (Kinoshita et al., 2013; Nakamura et al., 2000; Jeub et al., 2009; Ko et al., 2017) focus exclusively on reverberant speech, while others focus only on multi-speaker scenarios (Kraaij et al., 2005; Barker et al., 2017), environmental noise (Snyder et al., 2015; Reddy et al., 2020; Wichern et al., 2019; Hirsch & Pearce, 2000), or accented speech (Lander, 2022; Shi et al., 2021).
 Furthermore, certain types of digital perturbations, like special effects, computer-generated speech, and adversarial examples, which humans are usually invariant to (and thus ASR models are expected to be as well) are largely overlooked by existing benchmarks.
-To facilitate future research, SRB combines several benchmark datasets containing known noise sources to allow researchers to comprehensively evaluate the robustness of their models using a single benchmark.
-In the absence of comprehensive benchmarks, prior works have tried to evaluate ASR model robustness in various ways.
-For example, several works have used subsets of existing benchmarks to evaluate the model robustness [19, 20, 21], while others [22, 19, 58] have evaluated the robustness of models by computing their transcription accuracy on multiple natural speech datasets.
-A key shortcoming of these approaches is that the choice of benchmarks and datasets varies across studies, making it difficult to compare results.
-Recently, Huggingface Open ASR Leaderboard [23] has sought to mitigate this issue of inconsistent evaluation setups by developing a benchmark comprising several real-world speech datasets.
-Although evaluating models only on naturally noisy data may accurately reflect their real-world performance, it may not inform about the specific type of perturbations the models are weak against, because noise sources present in these datasets are not controlled or even fully known.
-For example, the crowdsourced recordings in Common Voice [24] contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers.
+SRB evaluates robustness under a wide range of scenarios, including those represented in prior works, and novel ones that have not yet received due attention, thus enabling comprehensive robustness evaluations via a single benchmark.
+Secondly, since there are several robustness benchmarks that researchers can choose from, the onus is on model developers to collect all relevant benchmarks and comprehensively evaluate their model.
+This has often resulted in model developers evaluating their models on disparate benchmarks  (Radford et al., 2023; Wen et al., 2016; Chen et al., 2022; Likhomanenko et al., 2020), which makes it hard to reliably compare performance and robustness across models.
+For example, several works have used subsets of existing benchmarks to evaluate the model robustness (Radford et al., 2023; Wen et al., 2016; Chen et al., 2022), or have evaluated the robustness of models by computing their transcription accuracy on multiple natural speech datasets (Likhomanenko et al., 2020; Radford et al., 2023; Hsu et al., 2021b).
+SRB covers a wide range of challenging ASR scenarios in a single benchmark, it obviates the need to select from various benchmarks, and, thus, can unify robustness evaluations across studies and make them comparable.
+Finally, robustness benchmarks that try to mitigate the above two shortcomings are often too coarse to reveal the specific scenarios and corruptions the model(s) struggle against.
+For example, Huggingface Open ASR Leaderboard(OAL, Srivastav et al.
+2023) evaluates models on several real-world speech datasets, which may accurately reflect their average-case real-world performance, it may not inform about the specific type of perturbations the models are weak against.
+This is because noise sources present in these datasets are not controlled or even fully known.
+For example, the crowdsourced recordings in Common Voice (Ardila et al., 2019) contain a variety of distortions including sensor noise from low-quality equipment, background noise, and mispronunciation by non-native speakers.
 If a model struggles with Common Voice, identifying the specific types of noise it is sensitive to is not a straightforward task.
-By evaluating models on a diverse set of known perturbations and noise sources, including adversarial attacks, SRB complements and extends benchmarks, like [23], that measure average case real-world performance and, thus, allows the robustness of ASR models to be comprehensively evaluated.
+In the design of SRB, we have ensured that the sources of noise (and variance, in general) in each utterance are limited and known, which allows users to pinpoint specific scenarios and/or perturbations under which their models struggle (e.g.
+Fig.
+3 and § 4).
 3 Speech Robust Bench
 Figure 1: An illustration of the processes involved in using our benchmark to evaluate the robustness of ASR models.
-At a high level, Speech Robust Bench (SRB) consists of four components: (1) Clean datasets, and a curated selection of inherently noisy datasets containing limited and known sources of noises and variations that are difficult to simulate; (2) Large bank of perturbations applicable on audio data; (3) Transcription extraction from ASR models and analysis; (4) Set of metrics for quantifying ASR model robustness, and facilitating comparisons among models.
-The basic workflow of SRB is the following (see Fig.
-1): first, recordings from the clean dataset(s) (§ 3.1) are perturbed by a large bank of perturbations (§ 3.2), resulting in several perturbed versions of each recording.
-Next, these perturbed recordings, the original clean recordings, and the recordings with inherent noise are transcribed using the target ASR model.
-The predicted and reference transcripts are compared, and the accuracy and robustness of each model under each type of noise are captured with various metrics (§ 3.3).
-To account for the relative differences in the level of distortion introduced by different perturbations, we estimate speech quality scores using appropriate models and use them to normalize the metrics.
-3.1 Dataset
-We use clean speech from Librispeech [26] and TEDLIUM [27].
-Librispeech contains professional recordings of audio books, and thus, has properly enunciated and cleanly recorded speech.
-Meanwhile, TEDLIUM contains professional recordings of TED talks and provides a degree of lexical and phonetic diversity that may be lacking in Librispeech.
-We chose these datasets due to their popularity in the ASR literature, however, any dataset containing utterances and reference transcripts would also be compatible with SRB.
-In addition to clean speech, SRB also uses inherently noisy speech containing specific types of perturbations that are hard to simulate, but are likely to be encountered by ASR models in the real-world.
-Specifically, SRB uses accented speech from Common Voice [24], and inter-personal conversations and far-field recordings from CHiME-6 [7] and AMI [8].
-3.2 Perturbations
-The taxonomy of the perturbations used in our benchmark is presented in Fig.
-2.
-The perturbations are of two broad types: (1) non-adversarial and (2) adversarial.
-The non-adversarial perturbations fall under one of four categories: (1.1) Inter-personal communication gathered in social settings using near-field and far-field microphones.
-This setting is representative of the environments in which voice-enabled smart devices are expected to operate.
-The data for this category are drawn from the ChiME [7] and AMI [8] corpora.
-(1.2) Environmental effects, such as white noise, environmental noise, spatial acoustics (room impulse response and echo) (1.3) Digital augmentations produced by standard audio processing operations, and special effects.
-The audio processing operations included amplitude gain, resampling, lowpass, and highpass filtering.
-The special effects included bass gain, treble gain, tempo increase, tempo decrease, speed increase, speed decrease, pitch increase, pitch decrease, chorus, tremolo, and phaser.
-All special effects are applied using SoX.
-[^1] https://sourceforge.net/projects/sox/ (1.4) Speech Variations are represented by speech from non-American speakers, and speech generated by a TTS model [59].
-Since users from a variety of ethnic backgrounds may interact with ASR models, they must be robust to a variety of accents.
-With the rise of generative AI, robustness to synthetic speech will also be important.
-Figure 2: Taxonomy of perturbations currently used in SRB.
-The adversarial perturbations we use are untargeted adversarial attacks whose objective is to prevent the correct transcription of input speech.
-We use two types of adversarial attacks: (2.1) utterance-specific and (2.2) utterance-agnostic attacks.
-The utterance-specific attack optimizes a perturbation for a given speech recording, such that a given model maximally mistranscribes.
-The utterance-agnostic attack works similarly but the perturbation is jointly optimized over many speech recordings.
-Thus, this perturbation can affect several speech recordings, even those that were not used during the optimization.
-This represents a more realistic scenario where a malicious actor tries to mount a denial-of-service style attack against an ASR model by introducing utterance-agnostic perturbation at some point in the transcription pipeline.
-Most perturbations, except for those categorized as informal speech, or speech variations, are applied at 4 levels of severity, which are modulated by adjusting the perturbation parameters.
-Details about the perturbations and their parameters are presented in Appendix A.
+Speech Robust Bench (SRB) evaluates the robustness of ASR models by a three-step process consisting of (1) scenario simulation, (2) transcription, and (3) metrics computation, as shown in Fig.
+1: first, various challenging speech recognition scenarios are simulated by applying a large bank of synthetic perturbations to clean speech datasets (§ 3.1), as well as by using inherently noisy speech datasets with limited and known sources of real noise and variations that are difficult to simulate.
+Next, the perturbed recordings, the original clean recordings, and the recordings with inherent noise are transcribed using the target ASR model.
+Finally, the predicted and reference transcripts are compared, and the accuracy and robustness of each model in each setting is captured with various metrics (§ 3.2).
+To account for the differences in the level of difficulty between scenarios, we also estimate speech quality scores using appropriate models and use them to calculate normalized metrics.
+Figure 2: Taxonomy of scenarios currently represented in SRB. Scenarios in dashed boxes have real-world recordings, while scenarios in solid boxes are simulated by digitally adding perturbations.
+3.1 Scenario Simulation
+The various speech recognition scenarios simulated by SRB are taxonomized in Fig.
+2, and can be divided into six high-level categories, namely (1) clean speech, (2) social gatherings, (3) speech variations, (4) environmental effects, (5) digital augmentations, and (6) adversarial attacks.
+The scenarios are described briefly below, while more details are given in Appendix A.
+(1) Clean speech: SRB uses clean speech for two purposes: to benchmark the baseline accuracy of ASR models, and to simulate various challenging scenarios by perturbing it.
+Clean speech is drawn from Librispeech (Panayotov et al., 2015) test-clean, TEDLIUM (Hernandez et al., 2018) release 3 test and MultiLingual LibriSpeech (MLS) (Pratap et al., 2020) test.
+LibriSpeech contains professional recordings of English audio books.
+Meanwhile, TEDLIUM contains professional recordings of English TED talks and provides lexical and phonetic diversity which LibriSpeech may lack.
+To increase the applicability of SRB to non-English and multi-lingual models we also include Spanish speech from MLS, which contains professionally recorded audio books in several languages.
+(2) Social Gatherings: The ability to transcribe speech from semi-formal or informal settings, as well as far-field audio is useful for models used in meeting rooms, smart homes, and even subtitle generation, thus SRB includes English speech from dinner parties and meetings recorded by (2.1) near- and (2.2) far-field mics from CHiME-6 (Barker et al., 2017) and AMI (Kraaij et al., 2005).
+(3) Speech Variations: ASR models must remain accurate under variations in pronunciation and prosody to serve diverse speakers.
+We therefore include (3.1) clean accented speech [^1] Accent annotations (excluding US English), and a DNSMOS score $\geq 3.4$ (Reddy et al., 2020).
+from English and Spanish subsets of Common Voice 17 (CV17, Ardila et al.
+2019) in SRB.
+To provide additional prosodic variability and to represent the increasing pervasiveness of generative AI, we also include synthetic speech generated by YourTTS (Casanova et al., 2022) (English) and Bark Suno (b) (Spanish) from transcripts of the three clean datasets from scenario (1) in the voices of English and Spanish speakers from VCTK (Yamagishi et al., 2019) and Bark Speaker Library (v2, Suno a), respectively.
+The following scenarios involve synthetic perturbation of all three clean datasets from scenario (1).
+(4) Environmental Effects: While noisy real speech datasets like CommonVoice (Ardila et al., 2019) and Switchboard (Godfrey et al., 1992) exist, the noise in them is not controlled or even known.
+Thus in SRB, we perturb clean speech to simulate (4.1) environmental noise, and (4.2) spatial acoustics.
+Concretely, we add real environmental noise from ESC-50 (Piczak, 2015), MS-SNSD (Reddy et al., 2019), MUSAN (Snyder et al., 2015) and WHAM!
+(Wichern et al., 2019) at Signal-to-Noise Ratios (SNR) of 10, 20, 30 and 40 dB.
+To simulate spatial acoustics, we add echo via SoX [^2] Available from https://sourceforge.net/projects/sox/.
+and simulate Room Impulse Response (RIR) via convolution with real and simulated RIRs from Ko et al.
+(2017).
+(5) Digital Augmentations: Digital media often undergoes processing and contains special effects, which are therefore included in SRB.
+Specifically, we include standard audio processing operations like amplitude gain, resampling, lowpass, and highpass filtering, (2.2) special effects like bass gain, treble gain, tempo increase, tempo decrease, speed increase, speed decrease, pitch increase, pitch decrease, chorus, tremolo, and phaser, and (2.3) Gaussian white noise.
+(6) Adversarial Attacks: Models used in high-stakes settings are prime targets for adversaries and thus must resist attempts to compromise their accuracy.
+We use two types of adversarial attacks in SRB: (2.1) utterance-specific and (2.2) utterance-agnostic attacks.
+The utterance-specific attack searches for a perturbation $\delta$, for a given speech recording $x$, such that a given model maximally mistranscribes it.
+To find $\delta$, we follow Madry et al.
+(2018) and use projected gradient descent to solve $\max_{\delta:\text{SNR}(\delta,x)\leq\epsilon}\mathcal{L}(M(x),y^{*})$, where $\mathcal{L}$ is a differentiable loss function, like CTC-Loss, between the model’s output $M(x)$ and the true transcript $y^{*}$, with $\epsilon\in[10,40]$.
+The utterance-agnostic attack is similar to the utterance-specific attack, except $\delta$ is optimized over a held-out set, $\mathcal{X}^{\text{dev}}$, instead of each test utterance.
+This represents a more realistic scenario where an attacker tries to mount a denial-of-service attack against an ASR model by introducing utterance-agnostic perturbation at some point in the transcription pipeline.
+We use the method of Neekhara et al.
+(2019) to find $\delta:\mathbbm{E}_{x\in\mathcal{X}^{\text{dev}}}\text{SNR}(\delta,x)\leq\epsilon\in[10,40]$ such that $CER(\{x+\delta|x\in\mathcal{X}^{\text{dev}}\})>\tau$ (see Alg.
+1), where $\mathcal{X}^{\text{dev}}$ is the dev split of LibriSpeech, TEDLIUM and MLS.
+Note of Extensibility and Usage: We have released our source code with instructions for reconstructing the data in SRB, and reproducing the results of this paper ( § 1).
+SRB can easily be extended to other languages and speech datasets using the provided scripts for extracting accented speech from any language in CV17, and for simulating scenarios 3.2-6 on any speech recording or dataset.
+We have also publicly released the data for all the above scenarios, except adversarial attacks and perturbed TEDLIUM recordings, on Huggingface Hub [URL omitted for double-blind].
+We made these exceptions because TEDLIUM’s license prohibits the distribution of derivatives, and the adversarial attacks must be computed separately for each target ASR model.
+To the extent possible, we encourage users to evaluate their models on the publicly released data to ensure reproducibility.
 3.2 Metrics
@@
 WER is computed as the word-level edit distance between the reference and the predicted transcripts, normalized by the length of the reference (see Appendix B for formal definitions).
-To measure the robustness of the model to input perturbations, we use WER Degradation (WERD), computed as $WER(\mathcal{X}_{p})-WER(\mathcal{X})$, where $\mathcal{X}$ contains clean audio, while the audio in $\mathcal{X}_{p}$ is perturbed by perturbation $p$.
-If $p$ is synthetically applied, both $\mathcal{X}$ and $\mathcal{X}_{p}$ contain the same utterances, otherwise $\mathcal{X}$ contains clean audios from LibriSpeech and TEDLIUM.
-When aggregating metrics (WER/WERD) over multiple perturbations, we follow the practice of [1] and divide the metric by a measure of difficulty, i.e., by the (estimated) speech quality degradation.
-This adds weight to errors on “easy” corruptions (less quality degradation) and under weighs errors on “harder” corruptions (more quality degradation) when computing averages.
+To measure the robustness of the model under challenging scenarios, we use WER Degradation (WERD), computed as $WER(\mathcal{X}_{s})-WER(\mathcal{X})$, where $\mathcal{X}$ and $\mathcal{X}_{s}$ are datasets containing clean speech and speech from scenario $s$, respectively.
+For scenarios (1)-(3.1) (see § 3.1), $\mathcal{X}_{s}$ is an inherently noisy dataset, and $\mathcal{X}$ will be LibriSpeech for English and Multi-Lingual LibriSpeech for Spanish.
+For scenarios (3.2)-(6), $\mathcal{X}$ is a clean dataset, and $\mathcal{X}_{s}$ is a perturbed version of $\mathcal{X}$.
+When aggregating metrics (WER/WERD) over multiple scenarios, we follow the practice of Hendrycks & Dietterich (2019) and divide the metric by a measure of difficulty, i.e., by the (estimated) speech quality degradation.
+This adds weight to errors on “easy” scenarios (less quality degradation) and underweights errors on “harder” scenarios (more quality degradation) when computing averages.
 We refer to the difficulty normalized versions of WER/WERD as Normalized WER/WERD (NWER/NWERD).
-We estimate speech quality using DNSMOS [60] and PESQ[61], which are models of human judgments of speech quality and predict Mean Opinion Scores (MOS) [62].
+We estimate speech quality using DNSMOS (Reddy et al., 2019) and PESQ(Rix et al., 2001; Miao Wang & ananda seelan, 2022), which are models of human judgments of speech quality and predict Mean Opinion Scores (MOS, Rec 2018).
 PESQ uses various signal processing methods to predict MOS, while DNSMOS uses DNNs to do the same.
 To compute speech quality degradation we compute PESQ and DNSMOS for each clean and noisy recording multiplied by -1 (lower values indicate less degradation).
-Since we are only interested in the relative degradation between perturbations, we normalize the scores to have mean 50 and standard deviation 25.
+Since we are only interested in the relative degradation between scenarios, we normalize the scores to have mean 50 and standard deviation 25.
+Note on usage: We use NWERD for non-adversarial scenarios (1-5) but WERD for adversarial attacks because adversarial attacks are model-specific and thus DNSMOS/PESQ scores for adversarially perturbed audio will be different for each model, which will lead to a different normalization during NWERD computation and make comparisons difficult.
 4 Evaluation
-We evaluate several recent and popular ASR DNNs (§4.1) using SRB and analyze the results to uncover fine-grained differences in their robustness to non-adversarial and adversarial input perturbations.
+We evaluate several recent ASR DNNs (§4.1) using SRB and analyze the results to uncover fine-grained differences in their robustness in various challenging scenarios.
 We further extend our analysis by measuring ASR model robustness for various sub-groups, namely English speech and non-English (Spanish) speech, and male and female speakers.
-Prior works [63, 64] observe that there is a disparity in transcription accuracy between subgroups.
+Prior works (Liu et al., 2022; Veliche & Fung, 2023) observe that there is a disparity in transcription quality between subgroups.
 Our analysis augments these observations by showing that inter-group disparities in robustness may also exist, thus demonstrating the utility of SRB in the broader field of trustworthy AI.
 4.1 Models
-For English speech, we evaluate Whisper[19] variants: large-v2 (wsp-lg), medium (wsp-md), small (wsp-sm), base (wsp-bs), and tiny (wsp-tn, wsp-tn.en), Wav2Vec-2.0 [38] variants: base (w2v2-bs), large (w2v2-lg), self-trained large [65] (w2v2-lg-slf), and Robust Wav2Vec [22], HuBERT [39] variants: large (hubt-lg) and XL (hubt-xl), Nvidia Canary [25] (cnry-1b), MMS [28] (mms-1b), Speech-T5 [66] (spch-t5) and DeepSpeech [67] (ds).
-For Spanish speech, we evaluate on the mono-lingual Wav2Vec base Spanish [68] (w2v2-bs-es), Wav2Vec XLSR Spanish [69] (w2v2-lg-es), wsp-lg, wsp-bs, wsp-tn, and mms-1b.
-We used the Huggingface implementations for all the models except cnry-1b (https://huggingface.co/nvidia/canary-1b) and ds (https://github.com/SeanNaren/deepspeech.pytorch).
-More details about the models are in Table 3 in the Appendix.
-Figure 3: The accuracy and robustness of English (top) and Spanish (bottom) ASR models on clean and perturbed data.
-Accuracy is measured by WER of the models on clean speech (a & d).
-Robustness is measured by the NWERD on non-adversarially perturbed speech (b & e) and WERD on adversarially perturbed speech (c & f).
-The markers indicate the dataset in (a & c), and the perturbation category in (b & e).
-The x axes are in ascending order of the values on the y axes.
-4.2 Accuracy on Clean Data
-Figures 3(a) and 3(d) show the WER of the models on clean English and Spanish, respectively.
-We see that cnry-1b is the most accurate model on both languages, followed by w2v2-lg-slf on English and wsp-lg on Spanish.
-4.3 Robustness of English ASR Models to Non-Adversarial Perturbations
-Fig.
-3(b) presents the NWERD for the English ASR models under non-adversarial noises and perturbations.
-We observe that wsp-lg emerges as the most robust model, followed by wsp-md and cnry-1b.
-Despite having the lowest overall WERD, wsp-lg is outperformed by wsp-md and cnry-1b on white noise, environmental noise, and special effects (see Appendix D for more detailed results).
-The massively multilingual mms-1b has the highest NWERD in social scenarios and far-field recordings category, indicating that it is unsuited for applications like voice assistants and smart home set-ups.
-We also notice that Wav2Vec-2.0 based models also struggle with the “speech variations” indicating that they are not suitable when users have diverse accents.
-Surprisingly, spch-t5 is the least robust model, and performs even worse than the much smaller and older ds.
-Similarly, wsp-md outperforms the newer and larger cnry-1b on almost all corruption categories.
-Figure 4: WERD of cnry-1b, wsp-md and wsp-sm on perturbations in the spatial acoustics and special effects categories.
-Upon closer examination, we see that cnry-1b has higher NWERD than its peers (wsp-md and wsp-sm) for two categories of perturbations: spatial acoustics and special effects.
-Upon looking at the WERD on each perturbation within these categories (Fig.
-4) we find that cnry-1b is more sensitive than its peers to echo, real room impulse responses, and speed and pitch modifications.
-This analysis demonstrates that SRB can evaluate the robustness of ASR models at multiple ganularities and can pinpoint the weaknesses of a given model.
+For English, we evaluate Whisper (Radford et al., 2023) large-v2, base, medium, small, and tiny (wsp-{lg,bs,md,sm,tn}), Wav2Vec-2.0 (Baevski et al., 2020) base, large, self-trained large (Xu et al., 2021), and Robust Wav2Vec (Likhomanenko et al., 2020) (w2v2-{bs,lg,lg-slf,lg-rob}), HuBERT (Hsu et al., 2021a) large and XL (hubt-{lg,xl}), Nvidia Canary (NVIDIA, ) (cnry-1b), Nvidia Parakeet RNN-T and CTC (NVIDIA, ) with 0.6B and 1.1B parameters (prkt-rnnt-{0.6,1.1}b, prkt-ctc-{0.6,1.1}b), MMS (Pratap et al., 2020) (mms-1b), Speech-T5 (Ao et al., 2022) (spch-t5), DeepSpeech (Amodei et al., 2016) (ds), and Speechbrain (Ravanelli et al., 2024) models with Conformer encoders, and transformer and RNN-T decoders.
+For Spanish speech, we evaluate mono-lingual Wav2Vec base Spanish (Wang et al., 2021b) (w2v2-bs-es), Wav2Vec XLSR Spanish (Conneau et al., 2020) (w2v2-lg-es), wsp-{lg,bs,tn}, and mms-1b.
+We used the Huggingface implementations where available, except ds (https://github.com/SeanNaren/deepspeech.pytorch).
+More details about the models are in Table 5.
+clean accent audio proc noise (env) noise (white) sFX social (FF) social (NF) spatial synth speech AVG Adv (UA) Adv (US) AVG
+Lang Model (WER) (NWERD) (WERD)
+EN prkt-rnnt-1.1b 5.9 11.6 8.7 4.2 1.6 6.4 48.3 39.8 11.8 3.0 15.0 10.9 69.1 40.0
+cnry-1b 6.0 13.9 18.2 4.4 2.7 15.0 45.7 36.7 15.3 5.7 17.5 14.8 61.7 38.3
+prkt-ctc-1.1b 6.0 16.9 10.3 4.2 3.2 9.6 44.6 35.0 13.5 5.9 15.9 8.4 71.2 39.8
+w2v2-lg-slf 7.7 41.0 30.7 13.1 17.6 20.3 69.5 67.6 26.3 15.2 33.5 7.0 41.0 24.0
+wsp-md 7.9 12.4 27.8 2.8 3.3 3.5 41.8 35.5 5.2 6.1 15.4 3.7 56.6 30.1
+wsp-lg 8.0 11.2 12.7 3.1 2.9 2.8 40.9 34.9 4.5 4.4 13.0 6.2 53.6 29.9
+hubt-xl 8.4 38.9 29.1 15.5 16.2 20.8 69.5 71.2 26.4 13.5 33.5 13.9 36.8 25.3
+wsp-bs 9.6 30.1 88.8 8.7 9.6 22.5 63.9 47.8 17.9 12.6 33.6 2.7 88.5 45.6
+w2v2-lg 9.7 60.6 39.5 19.0 26.0 24.1 77.3 79.9 37.3 17.7 42.4 16.6 31.1 23.8
+ES cnry-1b 3.2 699.9 21.7 17.4 7.6 30.5 - - 36.3 28.6 120.3 26.1 84.3 55.2
+wsp-lg 5.8 6.8 19.4 12.3 5.5 9.2 - - 5.5 24.6 11.9 13.7 65.0 39.4
+w2v2-lg-es 6.8 31.6 31.8 30.1 20.5 40.7 - - 89.0 104.1 49.7 33.9 71.0 52.4
+wsp-bs 14.8 62.0 133.2 43.1 25.8 60.4 - - 58.1 87.8 67.2 19.5 159.5 89.5
+$m$s-1b 15.7 26.3 27.9 32.7 9.3 43.3 - - 47.3 49.9 33.8 7.4 53.8 30.6
+w2v2-bs-es 25.7 45.6 55.9 44.9 25.0 58.4 - - 103.3 152.7 69.4 10.0 33.8 21.9
+Table 1: The utility and robustness of selected English and Spanish models (see Table LABEL:tab:full-results for more results).
+Utility is measured by WER of the models on clean speech.
+Robustness is measured by the NWERD on non-adversarially perturbed speech and WERD on adversarially perturbed speech.
+Adv (UA) refers to utterance agnostic attacks, while Adv (US) refers to utterance specific ones.
+The metrics are averaged over all datasets, perturbations, and severities in each category.
+4.2 Robustness of ASR Models
+Table 1 presents the utility and robustness of a subset of English and Spanish ASR models under non-adversarial and adversarial scenarios.
+The subset was selected to exclude small and/or less accurate models.
+The results of the excluded models, however, are used in § 4.3.
+4.2.1 Robustness in Non-Adversarial Scenarios
+English Models: In terms of average NWERD, we observe that wsp-lg emerges as the most robust model for non-adversarial scenarios, followed by prkt-rnnt-1.1b and wsp-md.
+Interestingly, cnry-1b, which is the top model on the Open ASR Leaderboard (OAL, Srivastav et al.
+2023), ranks 5th on SRB.
+This result highlights the fact that SRB provides a more rigorous assessment of a model’s robustness than existing benchmarks like OAL.
+We also see that SRB reveals subtle weaknesses and strengths of various models.
+For instance, we see that wsp-lg and wsp-md are significantly more robust to special effects (sFX) and spatial acoustics than other models, including cnry-1b.
+Figure 3: WERD of cnry-1b, wsp-md and wsp-lg on perturbations in the spatial acoustics and special effects categories.
+To identify the specific types of sFX and spatial acoustic perturbations against which cnry-1b lacks robustness, we plot the WERD on each perturbation within these categories (Fig.
+3) we find that cnry-1b is much more sensitive than its peers to echo, real room impulse responses, and speed and pitch modifications.
+This analysis demonstrates that SRB can evaluate the robustness of ASR models at multiple granularities and can pinpoint the weaknesses of a given model.
 Detailed results can be found in Figs.
-7 and 9 in the appendix.
-Takeaway: cnry-1b is the most accurate, but not most robust; Wav2Vec variants struggle against speech variations and social settings, thus, may not be suitable when users have diverse accents.
-4.4 Robustness of English ASR Models to Adversarial Perturbations
-Fig.
-3(c) shows the WERD of the models on utterance specific and utterance agnostic adversarial attacks.
-We observe that hubt-lg and hubt-xl emerge as the most robust models to utterance-specific adversarial attacks, followed by other models based on Wav2Vec-2.0 large, Whisper, and Canary.
-Interestingly, Wav2Vec models that exhibited mediocre robustness to non-adversarial perturbations are robust to utterance-specific attacks, while Whisper models and Canary, which were the most robust on non-adversarial perturbations, struggle on utterance specific attacks.
-We also note that most Wav2Vec models are considerably more robust to attacks against TEDLIUM than against LibriSpeech, and the opposite is true for whisper and Canary models.
-Under utterance-agnostic attacks, the best-performing models are mms-1b, wsp-bs, and wsp-sm.
+9 and 11 in the appendix.
+Given that no data augmentation, other than SpecAugment (Park et al., 2019), was used to train Whisper (Radford et al., 2023), this indicates that Whisper was trained on data that may have included digital media like music or movie soundtracks, and speech recorded in diverse acoustic environments – settings that may not be sufficiently represented in public data sources.
+Curating such diverse datasets is a promising direction for future work.
+We also note that despite being pre-trained on 60K hours of speech, Wav2Vec and Hubert models severely lack robustness.
+Particularly concerning is their weak performance on accented speech, social scenarios and spatial acoustics, which models are very likely to encounter in the real world.
+Takeaways: (1) Despite topping the Open ASR Leaderboard, cnry-1b is significantly less robust than wsp-lg, which is ranked 10 on OAL.
+cnry-1b particularly lacks robustness to special effects and spatial acoustics.
+(2) Wav2Vec variants struggle against accented speech and social settings, thus, may not be suitable when users have diverse accents.
+Spanish Models:We observe that wsp-lg is the most robust model against non-adversarial perturbations by some margin.
+We notice that all models, except wsp-lg, struggle against accented speech and yield high NWERS.
+cnry-1b is particularly, weak against accented speech with an NWERD of 700% (WERD=205%).
+Apart from accented speech, cnry-1b is quite robust on all other categories of non-adversarial perturbations.
+mms-1b is also fairly robust and, unlike other models, its NWERD does not vary erratically from one category to another.
+4.2.2 Robustness in Adversarial Scenarios
+English models: We observe that w2v2-lg achieves the lowest WERD and thus is the most robust model against utterance-specific adversarial attacks.
+Interestingly, while Wav2Vec models exhibited mediocre robustness to non-adversarial perturbations, they are more robust to utterance-specific attacks, than Whisper, Canary, and Parakeet, which were the most robust on non-adversarial perturbations.
+We also note from Fig.
+8 (in Appendix) that most Wav2Vec models are considerably more robust to attacks against TEDLIUM than against LibriSpeech, and the opposite is true for whisper and Canary models.
+Under utterance-agnostic attacks, the most robust models are mms-1b, wsp-bs, and wsp-sm.
 It is interesting to note that the smaller variants of Whisper limit the generalizability across utterances of the adversarial perturbations to a greater extent than their larger counterparts.
-Takeaway: Models that are most robust to non-adversarial perturbations, are mediocre against adversarial perturbations; cnry-1b is highly vulnerable to utterance specific attacks.
-4.5 Robustness vs. Model Attributes
+Takeaway: Wav2Vec models are most robust to adversarial attacks; Models that are most robust to non-adversarial perturbations, are mediocre against adversarial perturbations; Canary and Parakeet models are highly vulnerable to utterance specific attacks.
+Spanish models: On Spanish, w2v2-bs-es is the most adversarially robust model.
+Generally, we observe that Wav2Vec models exhibit better robustness than Whisper and Canary under both utterance-agnostic and utterance-specific perturbations.
+This is similar to the trends observed in English speech (Fig.
+8(c)).
+Detailed results can be found in Fig.
+10 in the appendix.
+Takeaway: General trends similar to English but WERD is higher when Spanish speech is attacked.
+4.3 Correlates of Robustness
+To glean insights that can inform future work, we have conducted the following analysis to model attributes that yield robust models.
+Specifically, we examine the impact of model size, architecture and accuracy, as well as training dataset size on robustness.
 To determine if the prevailing practice of training DNNs with more parameters on larger datasets is yielding improvements in robustness, we use robust linear regression to fit a line to WERD vs.
@@
 5(a) and 5(b), respectively.
-Increasing model size is correlated with improved robustness (lower WERD).To further isolate the impact of the model size we plot the NWERD of models from the same family in Fig.
-5(c), which have similar architectures and training datasets.
-We note that larger models are more robust in the Whisper and Wav2Vec-2.0 families, but, surprisingly, not in the HuBert family.
-In contrast, increasing training data appears to have only a minor influence on robustness (Fig.
-5(b)).
-Takeaway: Larger models tend to be more robust, while smaller models, even if they are trained on large datasets, are less robust. This runs somewhat counter to the prevailing wisdom [19, 22].
+Increasing model size is correlated with improved robustness (lower WERD).
+Figure 4: NWERD lineplot with non-adversarial and adversarial perturbations, three families of models.
+To further isolate the impact of the model size we plot the NWERD of models from the same family in Fig.
+4, which have similar architectures and training datasets.
+We note that larger models are more robust in the Whisper, Parakeet and Wav2Vec-2.0 families, but, surprisingly, not in the HuBert family.
+Next, we consider the model architectures.
+The architectures of the models used in this paper can be divided in to three categories: sequence-to-sequence (seq2seq) models like Whisper and Canary, encoder only models trained with CTC loss (Graves et al., 2006) like the Wav2Vec family, and RNN-T models which are capable of streaming such as some variants of Parakeet.
+From Fig.
+5(e) we see that in terms of non adversarial robustness RNN-T models outperform seq2seq and CTC models, but in terms of adversarial robustness CTC models achieve the lowest WERD.
+We also measure the robustness-utility trade-off by plotting WERD and NWERD for adversarial and non-adversarial perturbations, respectively, against WER on clean data in Figs.
+5(c) and 5(d).
+We observe that in both cases the relationship is positive, i.e.
+more accurate models tend to be more robust, however, the relationship between WERD on adversarial perturbations and clean WER is much weaker.
+Finally, we measure the impact of training data size and, find that increasing training data appears to have only a minor influence on robustness (Fig. 5(b)).
+Takeaway: (1) Larger models tend to be more robust, while smaller models, even if they are trained on large datasets, are less robust.
+This runs somewhat counter to the prevailing wisdom (Radford et al., 2023; Likhomanenko et al., 2020).
+(2) CTC models are more robust than seq2seq models to adversarial attack, but less robust than seq2seq and RNN-T models on non-adversarial perturbations.
+(3) Utility and robustness are positively correlated, but correlation is weaker for adversarial robustness.
 Figure 5: (a & b) WERD for all models with robust regression fitted line on non-adversarial (blue) and adversarial (orange) perturbations, plotted against (a) number of parameters, and (b) hours of training data.
-(c) NWERD lineplot with non-adversarial and adversarial perturbations, three families of models (color indicated).
-Models per family share similar architectures and training datasets.
-(d) Robustness of multi-lingual English (solid) and Spanish (hashed) models.
-4.6 Robustness for Population Sub-Groups
+(c % d) WERD and NWER on adversarial and non-adversarial perturbations are plotted against WER to illustrate the robustness-utility trade off.
+Pareto optimal points are highlighted.
+(e) Boxplot of WERD for models having various architectures.
+4.4 Disparity in Robustness Across Population Sub-Groups
 In the preceding analysis, we considered robustness aggregated over the entire population (i.e., dataset).
 However, populations are generally not homogeneous, and, thus, the robustness of the model may differ on various population sub-groups.
-Prior works have commonly analyzed sub-group fairness of ASR models by comparing the overall WER for each sub-group on a benchmark dataset [70].
+Prior works have commonly analyzed sub-group fairness of ASR models by comparing the overall WER for each sub-group on a benchmark dataset (Koenecke et al., 2020).
 It is possible that models that are fair on average, may not be fair under certain conditions.
@@
 We find that disparities indeed exist, with multi-lingual models generally being more robust for English than Spanish (Fig.
-5(d)), and most models being less robust for females than males.
-4.6.1 Robustness on Non-English (Spanish) Speech
-The clean accuracy of the Spanish ASR models is presented in Fig.
-3(d), and the NWERD under non-adversarial perturbations for each model is shown in Fig.
-3(e).
-We observe, that the two most accurate models, cnry-1b and wsp-lg, are also the two most robust models, but mms-1b, which had fairly high WER, is the third most robust model in terms of NWER.
-We also notice a general trend that all the models, except cnry-1b, struggle against “spatial acoustics” category, which is rather worrying since the influence of spatial acoustics is ubiquitous in the real-world.
-Considering adversarial attacks (Fig.
-3(f)), we observe that Wav2Vec models exhibit better robustness than Whisper and Canary under both utterance-agnostic and utterance-specific perturbations.
-This is similar to the trends observed in English speech (Fig.
-3(c)), and indicates that perhaps learned front-ends may be more adversarial robust than the spectrograms used by Whisper and Canary.
-Detailed results can be found in Fig.
-8 in the appendix.
-We also compare the robustness exhibited by multi-lingual models, wsp-lg, wsp-bs, cnry-1b and mms-1b on English and Spanish.
+6), and most models being less robust for females than males.
+4.4.1 Disparity in Robustness Across Languages in Multi-Lingual Models
+Figure 6: Comparing the robustness of multi-lingual on English (solid) and Spanish (hatched).
+We compare the robustness exhibited by multi-lingual models, wsp-lg, wsp-bs, cnry-1b and mms-1b on English and Spanish.
 The WERD of these models on both languages is presented in Fig.
-5(d).
+6.
 We observe that Whisper models achieve lower WERD on English speech than on Spanish on almost all perturbation categories, while cnry-1b and mms-1b achieve similar WERD on some categories.
@@
 Takeaway: Multilingual models are more robust on English than Spanish; cnry-1b and wsp-lg most robust on both languages; adversarial robustness results follow the same trend as English.
-4.6.2 Robustness and Gender
-To measure the disparity in prediction accuracy across genders (males/females), we compute the log of the ratio of the WERs of the ASR model on female and male speakers.
+4.4.2 Disparity in Robustness Across Genders
+To measure the disparity in transcription quality across genders (males/females), we compute the log of the ratio of the WERs of the ASR model on female and male speakers.
 We call this measure the Log WER Ratio (LWERR).
@@
 The LWERR for each dataset is shown in Fig.
-6.
+7.
 We note that, on average, the models are biased against females on LibriSpeech and Spanish Multilingual Librispeech (MLS-ES), and against males on TEDLIUM, Common Voice and AMI.
@@
 We propose SRB, a comprehensive benchmark designed to standardize the robustness evaluations of Automatic Speech Recognition (ASR) models.
-To the best of our knowledge, SRB is the first ASR robustness benchmark that contains long-form speech, and expands the set of distortions to include special effects, speaker attribute modification and adversarial attacks.
 We demonstrate the utility of SRB in evaluating the robustness of ASR models via several concrete examples, as well as its potential to facilitate evaluations of other aspects of trustworthy AI, like fairness.
@@
 universal adv 10.3 - - - 11.7 - - - 8.8 - - -
-voice conversion vctk 49.6 - - - 15.7 - - - 83.4 - - -
+synth. speech 49.6 - - - 15.7 - - - 83.4 - - -
 Table 3: Normalized DNSMOS and PESQ score for each perturbation.
+Name Subset Hours Utterances Speakers Male/Female
+LibriSpeech test-clean 5.4 2620 40 20/20
+TEDLIUM 3 test 3.76 1155 16 10/6
+Multi-Lingual LibriSpeech (es) test 10 2385 20 10/10
+CHiME-6 eval 5.25 13000 8 -
+AMI test 7.35 13168 16 8/8
+ESC-50 - 2.78 2000 - -
+MUSAN - 108.5 2016 - -
+WHAM! noise-test 9 3000 - -
+MS-SNSD noise-test 0.7 51 - -
+Table 4: Distributional statistics of speech (top) and noise (bottom) datasets used in SRB.
 Gaussian Noise:
@@
 Environmental Noise:
-We use the recordings of environmental noises from the test/eval subsets of ESC-50  [71], MS-SNSD [60], MUSAN [14] and WHAM [9].
+We use the recordings of environmental noises from the test/eval subsets of ESC-50  (Piczak, 2015), MS-SNSD (Reddy et al., 2019), MUSAN (Snyder et al., 2015) and WHAM (Wichern et al., 2019).
 We create a separate perturbed version of the clean data using each of these noise datasets.
@@
 Room Impulse Response:
-The simulated and real RIRs from [16] are applied to clean recordings.
+The simulated and real RIRs from (Ko et al., 2017) are applied to clean recordings.
 As a measure of intensity, RT60 is estimated for the simulated RIRs using Sabine’s formula with the room dimensions and absorption coefficient provided in the dataset.
-For the real RIRs, we compute the SRMR [72] using the implementation from https://github.com/aliutkus/speechmetrics/tree/master.
+For the real RIRs, we compute the SRMR (Santos & Falk, 2014) using the implementation from https://github.com/aliutkus/speechmetrics/tree/master.
 The severity is defined in increasing RT60s for the synthetic RIRs, and decreasing SRMR for the real RIRs.
@@
 Voice Conversion
-We use use YourTTS [59] from Coqui.ai [^2] https://github.com/coqui-ai/TTS to synthesize audio from textual transcripts in a given speaker’s style.
+We use use YourTTS (Casanova et al., 2022) from Coqui.ai [^3] https://github.com/coqui-ai/TTS to synthesize audio from textual transcripts in a given speaker’s style.
 The transcripts from the test clean subset of LibriSpeech are used.
-The target speakers are drawn from the VCTK corpus [73], which contains accented speech from 12 accents.
+The target speakers are drawn from the VCTK corpus (Yamagishi et al., 2019), which contains accented speech from 12 accents.
 For each transcript a random speaker is chosen to synthesize the audio.
-Untargeted Adversarial Attack:
-The adversarial perturbations are computed using an untargeted PGD adversarial attack.
+Crosstalk and Music
+We use crosstalk and music audios from MUSAN (Snyder et al., 2015). We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
+Accents
+We select a subset of audios from the test set of Common Voice 17.
+The selected audios satisfied the following criteria: (1) the speaker’s accent must be present in the metadata, (2) the accent must not be American, (3) should be clean.
+The last criterion is satisfied if the DNSMOS (Reddy et al., 2020) score of the recording is at least 3.4.
+The resulting subset contains 640 recordings.
+The most popular accent in this set is South Asian (India, Pakistan, Sri Lanka) (2̃5%), followed by British English (2̃5%).
+Inter-Personal Communications
+We CHiME-6 (Barker et al., 2017) and AMI (Kraaij et al., 2005) to obtain recordings of people in social scenarios (dinner party and meetings).
+In both these datasets, the speakers are recorded through lapel microphone and a room microphone resulting in near and far field recordings.
+We use both types of recordings and show separate results for them.
+We remove recordings that contain less than three words since they are often fillers.
+Uttarance Specific Adversarial Attack:
+The utterance specific adversarial perturbations are computed using the untargeted PGD adversarial attack implemented in robust_speech package (Olivier & Raj, 2022).
 The attack is computed as follows.
-First, the maximum possible L2 norm of the noise is determined by solving the equation for SNR for the norm of the noise.
+First, the maximum possible L2 norm of the noise is determined by solving the equation for SNR for the norm of the noise as follows.
 […]
@@
 where $\delta$ is the noise, $x$ is the audio signal and SNR is the upper bound on the SNR in the final signal.
-Then, we follow the approach of [43] and optimize $\delta$ using Projected Gradient Descent (PGD) to maximize the divergence between the true and predicted transcriptions.
-A.1 Crosstalk and Music
-We use crosstalk and music audios from MUSAN [14]. We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
-A.2 Accents
-We select a subset of audios from the test set of Common Voice 17.
-The selected audios satisfied the following criteria: (1) the speaker’s accent must be present in the metadata, (2) the accent must not be American, (3) should be clean.
-The last criterion is satisfied if the DNSMOS [10] score of the recording is at least 3.4.
-The resulting subset contains 640 recordings.
-The most popular accent in this set is South Asian (India, Pakistan, Sri Lanka) (2̃5%), followed by British English (2̃5%).
-A.3 Inter-Personal Communications
-We CHiME-6 [7] and AMI [8] to obtain recordings of people in social scenarios (dinner party and meetings).
-In both these datasets, the speakers are recorded through lapel microphone and a room microphone resulting in near and far field recordings.
-We use both types of recordings and show separate results for them.
-We remove recordings that contain less than three words since they are often fillers.
-Untargeted Utterance Agnostic Adversarial Attack:
-We use the method of [52] to compute utterance agnostic adversarial perturbations.
-Concretely, we compute a single perturbation vector that can cause the accuracy of the model on several (ideally all) inputs to degrade severely.
+Then, we follow the approach of (Madry et al., 2018) and optimize $\delta$ using Projected Gradient Descent (PGD) to maximize the divergence between the true and predicted transcriptions.
+Formally stated, the attack performs the following optimization:
+$\delta=\max_{\hat{\delta}:\left\lVert\hat{\delta}\right\rVert_{2}\leq\epsilon_{\text{SNR}}}L_{M}(x+\hat{\delta},y),$ (3)
+where $L_{M}$ is the loss function used to train the ASR model, $M$, such as CTCLoss or NLLLoss.
+Utterance Agnostic Adversarial Attack:
+We use the method of (Neekhara et al., 2019), as implemented in robust_speech package (Olivier & Raj, 2022), to compute utterance agnostic adversarial perturbations.
 The main difference between the universal attack and the PGD attack is that the latter computes a perturbation vector for each input, whereas the former computes a single perturbation that is expected to successfully attack any input to the model.
-Once we compute the perturbation we add it to the audio signal at the specified SNR.
-We use the robust_speech package [74] to generate adversarial perturbations.
+Formally, given a ASR model, $M$, and a development speech dataset, $\mathcal{X}^{\text{dev}}$ let $\mathcal{X}^{\text{dev}}_{\delta}=\{x+\delta|x\in\mathcal{X}^{\text{dev}}\}$ be the same dataset under additive perturbation $\delta$, and let $M(\mathcal{X}^{\text{dev}})$ and $M(\mathcal{X}^{\text{dev}}_{\delta}$ be the transcripts predicted by $M$ for $\mathcal{X}^{\text{dev}}$ and $\mathcal{X}^{\text{dev}}_{\delta}$.
+The utterance agnostic attack uses PGD to optimize $\delta$ such that $\left\lVert\delta\right\rVert_{\infty}\leq\epsilon$ and the Character-Error Rate (CER) (see § B) between $M(\mathcal{X}^{\text{dev}})$ and $M(\mathcal{X}^{\text{dev}}_{\delta})$ is at least $t$, i.e.
+$\text{CER}^{M}(\mathcal{X}^{\text{dev}}_{\delta},M(\mathcal{X}^{\text{dev}}))\geq t$ (using the notation from § B).
+Similar to the utterance-specific attack, the value of $\epsilon$ is determined by the maximum allowable SNR using eq.(2), except that $\ell_{\infty}$ norms are used instead of $\ell_{2}$ norms.
+The full algorithm is described in Algorith 1.
+Algorithm 1 Utterance Agnostic Attack Algorithm
+Speech data $\mathcal{X}^{\text{dev}}$ and , ASR model $M$ and its loss function, $L_{M}$ (e.g.
+CTCLoss, NLLLoss, etc.), allowed SNR $s$, learning rate, $\alpha$, max epochs $e_{max}$, max iterations per sample $i_{max}$, target attack success rate, $t_{sr}$, target Character-Error Rate (CER) (see § B), $t_{cer}$
+procedure CER($a,b$)
+  return $\text{EditDistance}(a,b)/len(b)$ $\triangleright$ $a$ and $b$ are character sequences
+end procedure
+procedure SuccessRate($\mathcal{X}$)
+  return $\sum_{x\in\mathcal{X}}I[CER(M(x+v),M(x))>t_{cer}]$
+end procedure
+procedure SNRToNorm($x$, SNR)
 […]
+end procedure
+$\epsilon\leftarrow\sum_{x\in\mathcal{X}^{\text{dev}}}SNRToNorm(x,s)/|\mathcal{X}^{\text{dev}}|$
+$v\leftarrow 0$
+$e\leftarrow 0$
+while $SuccessRate<t_{sr}$ and $e<e_{max}$ do
+  for $(x,y)\in\mathcal{X}^{\text{dev}}$ do $\triangleright$ $x$ is the audio, $y$ is the transcript
+   $i\leftarrow 0$
+   $r\leftarrow 0$
+   while $CER(M(x+v+r),M(x))>t_{cer}$ and $i<i_{max}$ do
+     $\Delta_{r}\leftarrow\alpha sign(\nabla_{r}0.5\left\lVert r\right\rVert_{2}-L_{M}(x+v+r,y))$
+     $r\leftarrow clip_{\epsilon}\{r-\Delta_{r}+v\}-v$
+     $i\leftarrow i+1$
+   end while
+   $v\leftarrow clip_{\epsilon}\{r+v\}$
+  end for
+  $e\leftarrow e+1$
+end while
+Once we compute the perturbation we add it to the test audios ($\mathcal{X}^{\text{test}}$) at the specified SNR using torchaudio.function.add_noise.
+For LibriSpeech, we use 500 utterances from test-dev as $\mathcal{X}^{\text{dev}}$ and test-clean as $\mathcal{X}^{\text{dev}}$.
+For TEDLIUM, we use the full dev and test sets as $\mathcal{X}^{\text{dev}}$ and $\mathcal{X}^{\text{test}}$.
+For Multi-Lingual LibriSpeech, we use 500 utterances from the dev set in the relevant language as $\mathcal{X}^{\text{dev}}$ and the full test set of the same language as $\mathcal{X}^{\text{test}}$.
 Appendix B Additional definitions
@@
 where $ED$ computes the edit distance.
+The Character Error Rate (CER) can also be defined similarly, by using the character-level edit distance in the above equation.
 For quantifying differences between (binary) genders, we measure the disparity in prediction accuracy across males and females by the Log WER Ratio (LWERR). Formally,
@@
 Most of the state of the art quantifies fairness in terms of predictive performance disparities.
-This occurs in the domain of fairness in ASR systems [63, 64, 70] and similar prevalence is found in other domains [75, 76].
-However, previous work in the domain of ASR also considered robustness disparities as an alternative fairness notion [77].
+This occurs in the domain of fairness in ASR systems (Liu et al., 2022; Veliche & Fung, 2023; Koenecke et al., 2020) and similar prevalence is found in other domains (Julia Angwin, ; Solans Noguero et al., 2023).
+However, previous work in the domain of ASR also considered robustness disparities as an alternative fairness notion (Nanda et al., 2021).
 Moreover, we argue that considering the dimension of robustness could give better sense of the expected disparities that could be observed when deployed in the wild, in the presence of diverse noisy conditions.
@@
 The WER of multilingual models is presented as English/Spanish.
-# Params (M) Data (hrs) WER
-Language Model model (abrv)
-EN canary-1b [25] cnry-1b 1,000.0 85,000 6.0
-deepspeech[67] ds 86.0 960 26.5
-hubert-large-ls960-ft[39] hubt-lg 317.0 60,000 8.4
-hubert-xlarge-ls960-ft[39] hubt-xl 964.0 60,000 8.4
-mms-1b-fl102 [78] $m$s-1b 964.6 55,000 22.8
-speecht5 asr[66] spch-t5 154.6 960 22.1
-wav2vec2-base-960h [38] w2v2-bs 95.0 960 11.3
-wav2vec2-large-960h [38] w2v2-lg 317.0 960 9.7
-wav2vec2-large-960h-lv60-self [65] w2v2-lg-slf 317.0 60,000 7.7
-wav2vec2-large-robust-ft-libri-960h [58] w2v2-lg-rob 317.0 63,000 8.9
-whisper-base [19] wsp-bs 74.0 680,000 9.6
-whisper-base.en [19] wsp-bs.en 74.0 563,000 5.1
-whisper-large-v2 [19] wsp-lg 1,550.0 680,000 8.0
-whisper-medium [19] wsp-md 769.0 680,000 7.9
-whisper-medium.en [19] wsp-md.en 769.0 563,000 4.1
-whisper-small [19] wsp-sm 244.0 680,000 8.3
-whisper-small.en [19] wsp-sm.en 244.0 563,000 4.0
-whisper-tiny [19] wsp-tn 39.0 680,000 11.3
-whisper-tiny.en [19] wsp-tn.en 39.0 563,000 10.1
-ES canary-1b [25] cnry-1b 1,000.0 85,000 3.2
-mms-1b-fl102 [78] $m$s-1b 964.6 55,000 15.7
-wav2vec2-base-10k-voxpopuli-ft-es [68] w2v2-bs-es 94.4 10,116 25.7
-wav2vec2-large-xlsr-53-spanish [69] w2v2-lg-es 315.4 54,350 6.8
-whisper-base [19] wsp-bs 74.0 680,000 14.8
-whisper-large-v2 [19] wsp-lg 1,550.0 680,000 5.8
-whisper-tiny [19] wsp-tn 39.0 680,000 23.3
+Lang Model Abrv. Params (M) Data (hrs) WER
+EN canary-1b (NVIDIA, ) cnry-1b 1,000.0 85,000 6.0
+deepspeech(Amodei et al., 2016) ds 86.0 960 26.5
+hubert-large-ls960-ft(Hsu et al., 2021a) hubt-lg 317.0 60,000 8.4
+hubert-xlarge-ls960-ft(Hsu et al., 2021a) hubt-xl 964.0 60,000 8.4
+mms-1b-fl102 (Pratap et al., 2023) $m$s-1b 964.6 55,000 22.8
+speecht5 asr(Ao et al., 2022) spch-t5 154.6 960 22.1
+wav2vec2-base-960h (Baevski et al., 2020) w2v2-bs 95.0 960 11.3
+wav2vec2-large-960h (Baevski et al., 2020) w2v2-lg 317.0 960 9.7
+wav2vec2-large-960h-lv60-self (Xu et al., 2021) w2v2-lg-slf 317.0 60,000 7.7
+wav2vec2-large-robust-ft-libri-960h (Hsu et al., 2021b) w2v2-lg-rob 317.0 63,000 8.9
+whisper-base (Radford et al., 2023) wsp-bs 74.0 680,000 9.6
+whisper-base.en (Radford et al., 2023) wsp-bs.en 74.0 563,000 5.1
+whisper-large-v2 (Radford et al., 2023) wsp-lg 1,550.0 680,000 8.0
+whisper-medium (Radford et al., 2023) wsp-md 769.0 680,000 7.9
+whisper-medium.en (Radford et al., 2023) wsp-md.en 769.0 563,000 4.1
+whisper-small (Radford et al., 2023) wsp-sm 244.0 680,000 8.3
+whisper-small.en (Radford et al., 2023) wsp-sm.en 244.0 563,000 4.0
+whisper-tiny (Radford et al., 2023) wsp-tn 39.0 680,000 11.3
+whisper-tiny.en (Radford et al., 2023) wsp-tn.en 39.0 563,000 10.1
+ES canary-1b (NVIDIA, ) cnry-1b 1,000.0 85,000 3.2
+mms-1b-fl102 (Pratap et al., 2023) $m$s-1b 964.6 55,000 15.7
+wav2vec2-base-10k-voxpopuli-ft-es (Wang et al., 2021b) w2v2-bs-es 94.4 10,116 25.7
+wav2vec2-large-xlsr-53-spanish (Conneau et al., 2020) w2v2-lg-es 315.4 54,350 6.8
+whisper-base (Radford et al., 2023) wsp-bs 74.0 680,000 14.8
+whisper-large-v2 (Radford et al., 2023) wsp-lg 1,550.0 680,000 5.8
+whisper-tiny (Radford et al., 2023) wsp-tn 39.0 680,000 23.3
 Table 5: Models used in our evaluations.
@@
 These figures may be referenced by the main text but were not included in the main body in the interest of space.
-Figures 7, and 8 present the breakdown by perturbation of the robustness of models on English and Spanish, respectively.
+Figure 8 provides an overview of the accuracy and robustness of the various models.
+Figures 9, and 10 present the breakdown by perturbation of the robustness of models on English and Spanish, respectively.
 Figure 11 presents a breakdown of robustness by severity.
+Figure 8: The accuracy and robustness of English (top) and Spanish (bottom) ASR models on clean and perturbed data.
+Accuracy is measured by WER of the models on clean speech (a & d).
+Robustness is measured by the NWERD on non-adversarially perturbed speech (b & e) and WERD on adversarially perturbed speech (c & f).
+The markers indicate the dataset in (a & c), and the perturbation category in (b & e).
+The x axes are in ascending order of the values on the y axes.
 Figure 9: NWERD of English models on different augmentations, averaged over all severities.
@@
 Table 6: Licenses of each of the considered datasets in SRB
+Table 7: Accuracy and robustness of ASR models on all datasets
+category clean accent audio proc noise (env) noise (white) sFX social (FF) social (NF) spatial synth speech AVG Adv (UA) Adv (US) AVG
+lang dataset model (abrv) WER NWERD WERD
+du MLS-DU cnry-1b 4.0 - 15.5 5.3 6.0 12.4 - - 12.2 - 10.3 - - -
+wsp-lg 7.3 - 14.3 2.6 6.5 4.1 - - 5.3 - 6.6 - - -
+$m$s-1b 14.6 - 21.6 8.1 12.0 14.3 - - 18.9 - 15.0 - - -
+wsp-bs 20.4 - 93.8 14.0 26.3 34.5 - - 24.7 - 38.6 - - -
+en LibriSpeech prkt-rnnt-1.1b 1.6 - 5.1 2.6 1.8 3.7 - - 8.0 3.0 4.0 - - -
+cnry-1b 1.7 - 12.3 3.0 2.7 12.7 - - 9.3 4.2 7.4 7.7 61.7 34.7
+prkt-rnnt-0.6b 1.8 - 11.2 3.4 3.0 6.0 - - 9.4 3.8 6.1 - - -
+w2v2-lg-slf 1.8 - 16.9 9.2 17.1 12.4 - - 14.4 5.5 12.6 12.5 58.0 35.2
+prkt-ctc-0.6b 2.0 - 10.9 3.5 3.4 7.3 - - 10.4 4.7 6.7 - - -
+prkt-ctc-1.1b 2.0 - 6.2 3.0 3.3 6.4 - - 9.1 4.4 5.4 - - -
+hubt-xl 2.1 - 18.2 11.1 16.3 13.3 - - 14.7 5.5 13.2 26.7 54.4 40.5
+hubt-lg 2.1 - 11.6 9.0 9.1 12.3 - - 15.0 6.5 10.6 34.9 50.0 42.5
+sb-cnfmr 2.6 - 9.4 11.0 15.2 11.3 - - 19.0 7.4 12.2 - - -
+w2v2-lg-rob 2.6 - 18.1 10.2 15.1 15.4 - - 16.9 5.9 13.6 45.8 61.0 53.4
+sb-cnfmr-rnnt 2.7 - 15.7 9.8 17.2 9.6 - - 20.4 8.4 13.5 - - -
+w2v2-lg 3.0 - 24.7 14.4 28.2 15.8 - - 21.8 8.6 18.9 29.0 44.3 36.6
+w2v2-bs 3.7 - 30.6 18.9 36.4 20.3 - - 26.9 9.6 23.8 29.5 53.6 41.5
+wsp-md 3.9 - 19.1 2.4 3.3 2.1 - - 3.8 5.3 6.0 1.7 51.9 26.8
+wsp-lg 3.9 - 6.9 2.0 3.5 1.6 - - 3.0 4.0 3.5 3.4 43.6 23.5
+wsp-sm.en 4.0 - 34.9 6.9 - 6.0 - - 3.9 6.2 11.6 - - -
+wsp-md.en 4.1 - 16.0 4.9 - 2.8 - - 0.9 4.8 5.9 - - -
+wsp-sm 4.3 - 27.9 4.0 5.4 4.5 - - 6.5 7.0 9.2 8.3 71.1 39.7
+wsp-bs.en 5.1 - 48.0 12.4 - 17.9 - - 11.2 8.1 19.5 - - -
+wsp-bs 5.9 - 63.6 6.6 11.0 14.8 - - 14.5 8.5 19.8 3.6 96.7 50.2
+wsp-tn.en 6.4 - 57.2 8.6 14.0 25.8 - - 17.1 9.2 22.0 9.4 90.2 49.8
+spch-t5 7.2 - 32.6 49.8 26.2 27.3 - - 67.9 64.3 44.7 52.4 106.4 79.4
+sbcrdnn 7.2 - - 11.5 33.9 - - - - - 22.7 - - -
+wsp-tn 8.2 - 68.4 13.8 17.4 26.1 - - 21.8 9.9 26.2 22.5 - 22.5
+ds 15.1 - 37.9 26.2 28.9 32.0 - - 46.2 18.0 31.5 - 62.9 62.9
+$m$s-1b 15.4 - 16.3 5.5 7.0 11.1 - - 11.9 8.9 10.1 6.2 44.6 25.4
+TEDLIUM cnry-1b 10.2 - 15.3 3.1 2.9 10.6 - - 11.7 1.5 7.5 21.9 - 21.9
+wsp-md 12.0 - 22.9 1.7 3.5 3.1 - - 4.3 0.8 6.1 5.7 61.3 33.5
+wsp-lg 12.1 - 12.5 2.5 2.6 2.3 - - 4.0 0.4 4.0 9.0 63.6 36.3
+wsp-sm 12.3 - 32.0 2.3 4.9 6.1 - - 5.6 1.6 8.8 2.2 69.8 36.0
+wsp-bs 13.3 - 67.6 5.8 8.8 19.2 - - 8.7 4.1 19.0 1.7 80.3 41.0
+w2v2-lg-slf 13.5 - 27.6 9.5 19.3 17.2 - - 18.8 9.9 17.0 1.5 24.0 12.8
+wsp-tn.en 13.7 - 50.9 7.2 10.9 29.3 - - 12.9 2.8 19.0 51.7 90.7 71.2
+wsp-tn 14.4 - 59.7 11.5 15.8 29.9 - - 16.2 4.3 22.9 1.3 110.9 56.1
+hubt-lg 14.7 - 19.9 9.2 12.1 16.2 - - 17.5 9.1 14.0 0.7 13.9 7.3
+hubt-xl 14.7 - 24.6 11.1 17.3 16.9 - - 18.2 8.2 16.0 1.2 19.1 10.2
+w2v2-lg-rob 15.2 - 24.4 8.7 15.7 18.7 - - 19.4 9.2 16.0 1.0 17.9 9.4
+w2v2-lg 16.5 - 31.2 12.9 24.7 18.8 - - 25.5 9.2 20.4 4.2 17.9 11.0
+w2v2-bs 18.8 - 37.0 16.5 30.0 23.3 - - 32.6 9.5 24.8 7.1 19.9 13.5
+$m$s-1b 30.2 - 19.0 6.2 7.3 12.1 - - 11.5 0.0 9.4 0.0 53.1 26.6
+spch-t5 37.0 - 48.8 31.8 15.2 27.5 - - 48.1 21.7 32.2 31.8 90.6 61.2
+ds 38.0 - 40.1 16.1 24.3 30.0 - - 48.7 7.3 27.8 15.3 51.5 33.4
+ami cnry-1b - - - - - - 31.7 13.9 - - 22.8 - - -
+hubt-lg - - - - - - 51.0 27.7 - - 39.4 - - -
+hubt-xl - - - - - - 51.1 27.6 - - 39.4 - - -
+$m$s-1b - - - - - - 71.2 55.2 - - 63.2 - - -
+prkt-ctc-0.6b - - - - - - 34.5 14.5 - - 24.5 - - -
+prkt-ctc-1.1b - - - - - - 31.6 13.6 - - 22.6 - - -
+prkt-rnnt-0.6b - - - - - - 37.0 16.7 - - 26.9 - - -
+prkt-rnnt-1.1b - - - - - - 35.9 17.0 - - 26.4 - - -
+sb-cnfmr - - - - - - 63.0 35.4 - - 49.2 - - -
+sb-cnfmr-rnnt - - - - - - 58.6 31.6 - - 45.1 - - -
+sbcrdnn - - - - - - 91.2 - - - 91.2 - - -
+spch-t5 - - - - - - 304.3 221.9 - - 263.1 - - -
+w2v2-bs - - - - - - 66.0 32.9 - - 49.4 - - -
+w2v2-lg - - - - - - 59.6 30.8 - - 45.2 - - -
+w2v2-lg-rob - - - - - - 51.1 24.7 - - 37.9 - - -
+w2v2-lg-slf - - - - - - 50.6 26.2 - - 38.4 - - -
+wsp-bs - - - - - - 39.5 19.1 - - 29.3 - - -
+wsp-lg - - - - - - 29.3 15.9 - - 22.6 - - -
+wsp-md - - - - - - 29.2 15.8 - - 22.5 - - -
+wsp-sm - - - - - - 31.4 17.0 - - 24.2 - - -
+wsp-tn - - - - - - 48.9 21.2 - - 35.0 - - -
+wsp-tn.en - - - - - - 41.7 19.5 - - 30.6 - - -
+chime cnry-1b - - - - - - 55.6 28.8 - - 42.2 - - -
+ds - - - - - - 90.6 84.0 - - 87.3 - - -
+hubt-lg - - - - - - 82.1 55.2 - - 68.7 - - -
+hubt-xl - - - - - - 81.1 54.6 - - 67.8 - - -
+$m$s-1b - - - - - - 126.7 98.8 - - 112.7 - - -
+prkt-ctc-0.6b - - - - - - 56.6 27.4 - - 42.0 - - -
+prkt-ctc-1.1b - - - - - - 53.4 26.7 - - 40.0 - - -
+prkt-rnnt-0.6b - - - - - - 60.5 28.3 - - 44.4 - - -
+prkt-rnnt-1.1b - - - - - - 55.9 27.3 - - 41.6 - - -
+sb-cnfmr - - - - - - 92.5 74.2 - - 83.3 - - -
+sb-cnfmr-rnnt - - - - - - 84.3 60.5 - - 72.4 - - -
+sbcrdnn - - - - - - 92.9 87.3 - - 90.1 - - -
+spch-t5 - - - - - - 527.7 388.7 - - 458.2 - - -
+w2v2-bs - - - - - - 88.7 64.6 - - 76.7 - - -
+w2v2-lg - - - - - - 86.9 61.7 - - 74.3 - - -
+w2v2-lg-rob - - - - - - 83.4 52.3 - - 67.8 - - -
+w2v2-lg-slf - - - - - - 81.6 52.0 - - 66.8 - - -
+wsp-bs - - - - - - 83.4 35.5 - - 59.4 - - -
+wsp-lg - - - - - - 48.6 21.6 - - 35.1 - - -
+wsp-md - - - - - - 50.6 22.8 - - 36.7 - - -
+wsp-sm - - - - - - 59.2 24.9 - - 42.1 - - -
+wsp-tn - - - - - - 98.6 46.3 - - 72.4 - - -
+wsp-tn.en - - - - - - 80.4 35.2 - - 57.8 - - -
+CV cnry-1b - 4.6 - - - - - - - - 4.6 - - -
+ds - 93.4 - - - - - - - - 93.4 - - -
+hubt-lg - 14.1 - - - - - - - - 14.1 - - -
+hubt-xl - 12.9 - - - - - - - - 12.9 - - -
+$m$s-1b - 18.1 - - - - - - - - 18.1 - - -
+prkt-ctc-0.6b - 5.8 - - - - - - - - 5.8 - - -
+prkt-ctc-1.1b - 5.6 - - - - - - - - 5.6 - - -
+prkt-rnnt-0.6b - 5.1 - - - - - - - - 5.1 - - -
+prkt-rnnt-1.1b - 3.8 - - - - - - - - 3.8 - - -
+sb-cnfmr - 20.3 - - - - - - - - 20.3 - - -
+sb-cnfmr-rnnt - 19.1 - - - - - - - - 19.1 - - -
+sbcrdnn - 44.8 - - - - - - - - 44.8 - - -
+spch-t5 - 89.7 - - - - - - - - 89.7 - - -
+w2v2-bs - 25.4 - - - - - - - - 25.4 - - -
+w2v2-lg - 20.0 - - - - - - - - 20.0 - - -
+w2v2-lg-rob - 13.9 - - - - - - - - 13.9 - - -
+w2v2-lg-slf - 13.5 - - - - - - - - 13.5 - - -
+wsp-bs - 10.0 - - - - - - - - 10.0 - - -
+wsp-lg - 3.7 - - - - - - - - 3.7 - - -
+wsp-md - 4.1 - - - - - - - - 4.1 - - -
+wsp-sm - 6.5 - - - - - - - - 6.5 - - -
+wsp-tn - 13.9 - - - - - - - - 13.9 - - -
+wsp-tn.en - 12.7 - - - - - - - - 12.7 - - -
+es MLS-ES cnry-1b 3.2 - 16.4 9.6 8.2 16.0 - - 10.9 - 12.2 26.1 84.3 55.2
+wsp-lg 5.8 - 14.1 7.4 5.9 4.0 - - 2.0 - 6.7 13.7 65.0 39.4
+w2v2-lg-es 6.8 - 21.2 14.7 22.1 19.0 - - 26.3 - 20.6 33.9 71.0 52.4
+wsp-bs 14.8 - 90.9 22.7 27.3 32.5 - - 18.0 - 38.3 19.5 159.5 89.5
+$m$s-1b 15.7 - 18.5 19.7 37.1 20.0 - - 14.9 - 22.0 7.4 53.8 30.6
+wsp-tn 23.3 - 124.3 45.0 52.3 57.9 - - 41.8 - 64.3 43.1 269.9 156.5
+w2v2-bs-es 25.7 - 32.4 20.4 24.2 24.1 - - 32.4 - 26.7 10.0 33.8 21.9
+CV:es cnry-1b - 205.0 - - - - - - - - 205.0 - - -
+$m$s-1b - 7.7 - - - - - - - - 7.7 - - -
+w2v2-bs-es - 13.3 - - - - - - - - 13.3 - - -
+w2v2-lg-es - 9.3 - - - - - - - - 9.3 - - -
+wsp-bs - 18.2 - - - - - - - - 18.2 - - -
+wsp-lg - 2.0 - - - - - - - - 2.0 - - -
+wsp-tn - 30.6 - - - - - - - - 30.6 - - -
+fr MLS-FR cnry-1b 6.1 - 15.2 5.2 7.3 13.0 - - 10.0 - 10.1 - - -
+wsp-lg 7.7 - 15.5 3.5 8.3 5.6 - - 5.7 - 7.7 - - -
+$m$s-1b 23.6 - 21.0 6.9 12.5 12.9 - - 15.6 - 13.8 - - -
+wsp-bs 26.0 - 124.8 16.7 44.4 47.9 - - 25.9 - 51.9 - - -